### v1

# Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding

Abstract
Recent advances in vision-language models have significantly propelled video understanding. Existing datasets and tasks, however, have notable limitations. Most datasets are confined to short videos with limited events and narrow narratives. For example, datasets with instructional and egocentric videos often document the activities of one person in a single scene. Although some movie datasets offer richer content, they are often limited to short-term tasks, lack publicly available videos and frequently encounter data leakage given the use of movie forums and other resources in LLM training. To address the above limitations, we propose the Short Film Dataset (SFD) with 1,078 publicly available amateur movies, a wide variety of genres and minimal data leakage issues. SFD offers long-term story-oriented video tasks in the form of multiple-choice and open-ended question answering. Our extensive experiments emphasize the need for long-term reasoning to solve SFD tasks. Notably, we find strong signals in movie transcripts leading to the on-par performance of people and LLMs. We also show significantly lower performance of current models compared to people when using vision data alone.
1 Introduction
Recent advances in vision-language models [33, 31, 9, 10, 61, 60, 16, 15, 11, 68] show significant promise for enhancing machine perception. Despite their remarkable progress, current datasets are still constrained by some limitations. Most of them consist of short videos [20, 1, 13, 69, 72, 17, 55, 27, 32, 19, 65], often under one minute, and primarily target short-term tasks e.g., action recognition and video retrieval, often requiring only a few seconds of visual content for problem-solving [39, 28]. For long-form video understanding, the community has turned to three video categories. (i) Egocentric videos, an emerging domain featuring extended first-person sequences that capture a person’s continuous actions [18, 12, 24, 34, 39, 52, 51] or chronicle daily activities, offering tasks from object manipulation to social interactions [18]. Nevertheless, while such datasets excel in video duration, they lack narrative depth, focusing mainly on immediate visuals rather than rich storytelling; (ii) Instructional videos [41, 71], which depict a wide variety of human tasks and focus on the procedural aspect, are often brief and deficient in storytelling; (iii) Movies, unlike egocentric and instructional videos, provide complex plots and extended duration, making them an ideal testing ground for developing models that can comprehend long-form video narratives, see Figure 1.
Several movie datasets have been proposed in the literature [48, 4, 56, 29, 53, 21, 64, 22, 59, 66, 49, 40, 44, 7, 6, 38] offering rich narrative-driven content and various tasks. These tasks include clip retrieval [4], video question answering [56, 29, 38], audio description [53, 21], semantic role labeling [49], scene segmentation [22, 44], and visual scene recognition [6]. However, current movie datasets suffer from three main limitations. (i) Accessibility: these datasets are not publicly available as they typically contain commercial copyrighted movies (see Figure 2). (ii) Clip duration: they often consist of incomplete and short clips [56, 29, 45, 64] (i.e., less than 4 minutes, see Figure 2), limiting their potential for long-form understanding with evolving storylines. Moreover, these datasets frequently suffer from ill-designed benchmark questions with insufficient narrative focus [54]. (iii) Data Leakage: as these datasets comprise well-known commercial movies, modern Large Language Models (LLMs) and Vision-Language Models (VLMs) have likely been exposed to some form of movie information (i.e., synopses, reviews, discussions, subtitles, blog posts). Figure 3 showcases this, where given only movie titles, modern LLMs can obtain high accuracy when answering questions from main movie datasets. This data leakage leads to biased benchmarks and ineffective training.
To address the above limitations, we introduce the Short Film Dataset (SFD), a novel video benchmark comprising 1,078 short films, totalling over 243 hours of videos with an average duration of 13 minutes per film—significantly longer than existing datasets (see Figure 2). Unlike copyrighted films, SFD includes publicly accessible amateur films of diverse genres from YouTube. Although these films are shorter than traditional movies, they feature complex narratives that unfold through key sequences of events and multiple character interactions. Most importantly, with minimal exposure to LLMs, our dataset is less prone to data leakage issues plaguing other movie datasets (see Figure 3).
Aiming towards long-term story-level multimodal video understanding, we propose two question-answering tasks on SFD: (a) following recent trend [56, 45], we propose Multiple-Choice Question answering (MCQ); (b) We also introduce Open-Ended Question answering (OEQ), a more generative and unrestricted QA format that more realistically and objectively measures a model’s ability to comprehend long-form videos. All questions and answers are generated by LLMs from movie descriptions and undergo thorough manual curation, ensuring they accurately reflect the settings, characters, storylines, and themes of the movies.
To validate the advantages of SFD, we conduct extensive experiments and quantify data leakage, the benefits of long-term reasoning and comparison to human performance while using state-of-the-art vision and language models. Our findings are: (i) Unlike existing VideoQA datasets, SFD provides story-level tasks defined on public data with minimal data leakage (Section 3.1); (ii) our long-term reasoning experiments in Section 3.2 confirm the advantage of movie-level video understanding for solving SFD tasks; (iii) the evaluation in Section 3 reveals a large gap between recent VideoQA methods and human performance in vision-only and multimodal settings. SFD, hence, provides a unique benchmark for evaluating and advancing new methods for story-level long-term video understanding. Our dataset, code, and models are publicly available at https://shortfilmdataset.github.io.
2 Short Film Dataset
Short films are motion pictures that typically last between 5 and 20 minutes. As illustrated in Figure 4, short films span a variety of styles (e.g., narrative fiction, documentary, animation, etc.) and genres (e.g., action, drama, comedy, horror, etc.). Filmmakers often use the short film format to explore new ideas, techniques, and storytelling styles. Short films, hence, have more of an experimental rather than commercial value and are often made publicly available. Samples from the dataset can be found in Appendix H.
The amount of publicly available short films has significantly increased over the last years (see Figure 6(c)). We take advantage of this development and create a Short Film Dataset (SFD) to foster research on story-level video understanding. In the remaining part of this section, we describe the procedure for collecting videos and corresponding metadata (Section 2.1) as well as generating question-answer pairs (Section 2.2). Section 2.3 presents the analysis and statistics of SFD. The overall pipeline for SFD creation is illustrated in Figure 5.
2.1 Data Collection
To build our dataset, we capitalize on the abundance of short films on video-sharing platforms. We specifically target the Omeleto YouTube channel111https://www.youtube.com/@Omeleto and its subchannels, including @OmeletoComedy and @OmeletoDrama, among others. These channels promote high-quality and award-winning short films, some of which have been recognized by Oscars and BAFTA awards. We download videos, subtitles, and associated metadata from YouTube using yt-dlp222https://github.com/yt-dlp/yt-dlp.
As speech transcripts were originally available only for a portion of the dataset, we have used WhisperX [3] to extract missing transcripts. The metadata accompanying each short film includes the movie title, a logline which is a concise one-sentence summary of the plot, and additional movie details such as genre, release year, region of origin, and language. Each film also comes with a detailed description that offers a deeper understanding of various elements such as a synopsis providing a brief narrative overview, the director’s inspirations, storytelling approach, actors’ performances, and other external information. To maintain clarity and objectivity in our dataset, we retain only the synopsis and exclude other subjective or ambiguous elements of the description.
2.2 Question and Answer Generation
To promote the development of long-form and story-oriented video understanding, we complement SFD with two Video Question Answering (VideoQA) tasks: Multiple-Choice Question answering (MCQ) and Open-Ended Question answering (OEQ). We first automatically generate question-answer pairs by prompting GPT-4 [42] with movie titles, loglines and synopses. We then manually curate questions and answers to ensure their relevance and correctness. We describe the generation process in more details below.
Automatic QA pair generation. For each film we obtain human-written information, i.e. movie titles, loglines and synopses from Omeleto YouTube channels. We then prompt GPT-4 [42] to propose relevant question-answer pairs from movie information, with dedicated prompts to avoid common mistakes. In particular, we emphasise direct and clear questions, with brief and specific answers to avoid speculation, repetition and ambiguity. Our example prompts are available in Appendix I. This stage results in 7,104 question-answer (QA) pairs that we use for the open-ended question-answering task.
Automatic Distractor Generation. To construct the MCQ task, we use GPT-4 [42] to complement each question with 4 distractors. The distractors are crafted to provide plausible yet incorrect answers that fit within the movie’s context. To this end, our prompts are designed with specific criteria: distractors should be syntactically similar to the correct answer while being semantically different. Our distractors also incorporate diverse misdirections relative to movies (character confusion, plot adjustments), consider the movie genre and plausible scenarios, and include accurate but irrelevant information from other parts of the synopsis. This stage results in 7,104 Multiple-Choice Questions (MCQs), with 5 options per question with one correct and four incorrect answers.
Manual QA Curation. To ensure the quality of questions and answers, we perform a thorough manual curation based on the following criteria. We remove questions that cannot be answered based on the synopsis, including factually incorrect, ambiguous, or subjective ones. Additionally, questions considered overly simplistic or answerable through external knowledge are excluded. This stage results in 4,885 MCQs from 1,078 movies.
2.3 Dataset Analysis and Statistics
Our SFD dataset contains a total of 1,078 unique films ranging in duration from 5 to 37 minutes and having an average length of 13 minutes (see Figure 6(a)). Overall, these movies amount to 243 hours and cover distinct genres, as illustrated in Figure 6(d). Each movie in the dataset is accompanied by a title, a logline which averages around 15 words, and a synopsis of about 97 words, equivalent to approximately 2 sentences, as shown in Figure 6(b). We use this metadata for question and answer generation as described in Section 2.2. The majority of films in SFD are in English and mainly originate from North America and Europe (see Figures 6(f) and 6(e)). Examples of the dataset’s diversity can be seen in Figure 4. Additional data preprocessing and statistics are available in Appendix C, including clips, captions, location tags, and face tracks.
In terms of video understanding tasks, SFD comprises 4,885 MCQs and OEQs, averaging 4.53 questions per film. Table 1 shows the comparison to both general video understanding (videoQA) and movie-oriented (movieQA) question-answering datasets. Notably, SFD exhibits the longest average video duration (821 seconds) among all datasets. Compared to videoQA datasets, SFD holds the properties of multimodality and long-term nature, similar to most movieQA datasets. Compared to other movieQA ones, SFD offers advantages in public availability of videos, data leakage (i.e., not exposed to LLMs), and story completeness (using full movies instead of trimmed clips, potentially impacting narrative coherence).
Dataset | Venue | Annotation | Avg. | #QA Pairs | Multimodal | Long-Term | Accessible | Unknown | Full |
Length (s) | To LLMs | Movies | |||||||
General VideoQA datasets | |||||||||
MSRVTT-QA [67] (test) | ACM 2017 | Auto | 15 | 72,820 | ✗ | ✗ | ✓ | ✓ | - |
MSVD-QA [8] (test) | ACM 2017 | Auto | 10 | 13,156 | ✗ | ✗ | ✓ | ✓ | - |
TGIF-QA [23] (test) | CVPR 2017 | Auto | 3 | 25,751 | ✗ | ✗ | ✓ | ✓ | - |
ActivityNet-QA [75] (test) | AAAI 2019 | Manual | 180 | 8,000 | ✗ | ✗ | ✓ | ✓ | - |
How2QA [50] (test) | EMNLP 2020 | Manual | 60 | 4,400 | ✗ | ✗ | ✓ | ✓ | - |
NeXT-QA [65] (test) | CVPR 2021 | Manual | 44 | 9,178 | ✗ | ✗ | ✓ | ✓ | - |
iVQA [35] | ICCV 2021 | Manual | 18 | 10,000 | ✗ | ✗ | ✓ | ✓ | - |
EgoSchema [39] | NeurIPS 2023 | Manual + Auto | 180 | 5,000 | ✗ | ✓ | ✓ | ✓ | - |
MovieQA datasets | |||||||||
MovieQA [56] (test) | CVPR 2016 | Manual | 203 | 6,462 | ✓ | ✓ | ✗ | ✗ | ✗ |
TVQA [29] (test) | EMNLP 2018 | Manual | 76 | 15,253 | ✓ | ✗ | ✗ | ✓ | ✗ |
LVU [64] (test) | CVPR 2021 | Manual | 220 | 1,223 | ✓ | ✓ | ✓ | ✗ | ✗ |
MovieChat [54] (test) | CVPR 2024 | Manual | 459 | 2,417 | ✓ | ✓ | ✗ | ✓ | ✗ |
CinePile [45] (test) | arXiv 2024 | Manual + Auto | 160 | 4,940 | ✓ | ✓ | ✓ | ✓ | ✗ |
SFD (Ours) | Manual + Auto | 821 | 4,885 | ✓ | ✓ | ✓ | ✓ | ✓ |
Question Analysis.
The generated question-answer pairs are long and diverse: the median answer length is 10 words and the vocabulary contains 8,928 different words. Furthermore, to ensure that the generated QAs cover various aspects of movies, we categorize them into four types (see Table 2):
1. Setting-related questions focus on the location and time of the movie. They may pertain to geographical locations (e.g., a specific city or country), a type of place (e.g., home, office), a historical period (e.g., a specific era), or a time of day (e.g., morning, night).
2. Character-related questions focus on the individuals portrayed in the movie. They may concern personal information (age, gender, profession), traits, motivations, and relationships (siblings, friends). Note that character development, as part of the narrative, is excluded from this category.
3. Story-related questions focus on the key sequence of events and interactions that compose the plot and narrative arc of the movie.
4. Theme-related questions focus on the underlying message or central idea that the movie explores and conveys to its audience.
Question Type | Questions | Answers |
Setting | Where do John and Emma have their conversation? | On the bus. |
Which city is the backdrop? | Liverpool. | |
What decade does the movie take place in? | 1990s. | |
Character | Who is the main character? | Josh. |
What is the nationality of the soldier who becomes a prisoner? | German. | |
How is Mr. Jones described? | As stern and harsh. | |
Story | Can you name a specific tactic Tommy uses against Tiny? | He steals Tiny’s food at dinner and distracts her at a school-wide relay race. |
What is the imminent threat to Greg’s residence? | The house is in foreclosure and the bank is about to repossess it. | |
What triggers Carla to believe her husband Frank will surprise her for their anniversary? | A bouquet left on her doorstep. | |
Theme | What is a major theme? | The theme revolves around reconnecting and the complexities of honesty between exes. |
What main theme is explored through Sarah’s experience? | The theme of romantic expectations versus present realities. | |
What is the main theme? | The processing of grief and reconciliation between siblings. |
3 Experiments
3.1 Data Leakage
Modern LLMs, pre-trained on vast internet data, risk being exposed during training to information about commercial films, such as synopses, reviews, scripts, and transcripts. This issue may result in models recalling answers directly without analyzing the video content, leading to biased benchmarks. In this section, we quantitatively assess the extent of data leakage on datasets. Specifically, we prompt LLMs to answer open-ended questions using only the movie title and compute the accuracy of responses, following [37]. We experiment with 3 movie datasets: MovieQA [56], LVU [64] and our SFD, 5 open-sourced models: Gemma 2B [57], Mistral 7B/8x7B [25], LLaMA-3 8B/70B [58] and 4 commercial models Claude 3 Haiku/Sonnet, and GPT-3.5/4 [42]. Figure 3 reports the results of the data leakage experiment and further details can be found in Appendix G.
We observe that both MovieQA and LVU suffer from data leakage, reaching up to 71.3% and 76.0% accuracies. In contrast, thanks to the low presence of amateur film on the internet, SFD exhibits a maximum accuracy of 36.0%, indicating a low leakage issue. Furthermore, as expected, the extent of data leakage correlates with the knowledge level reflected by MMLU score [14] (see Appendix section G for more experiments), indicating that LLMs with more knowledge exhibit higher levels of memorization, leading to worse leakage issues. For instance, for both MovieQA and LVU, the zero-shot accuracy increases from approximately random level (20%) to more than 70% as the model size rises. Meanwhile, SFD maintains stable and low accuracies ranging between 19.7% and 36.0%, regardless of LLM knowledge variation, further indicating its low data leakage. This experiment reveals that relying solely on existing datasets to evaluate new methods is insufficient. Instead, our SFD offers a more objective and reliable test bed for long-term video understanding.
3.2 Temporal Window Study
To confirm that our tasks require long-form video understanding, we conduct a temporal window study with benchmarks at three levels: shot-, scene-, and movie-level. A shot, defined as a continuous video clip, typically amounts to around 151 per movie. At the Shot-Level, the model uses data from a single shot, including partial subtitles and visual data, with predictions aggregated by taking the maximum logits. Scene-Level inference aggregates data from approximately 10 shots in a similar manner. Movie-Level inference utilizes the entire film’s data, as detailed in Section 3.3. To further analyze the behaviors of each modality, we ablate at each temporal level three combinations of modalities: Vision, Language, and multimodal Vision-Language. This study is conducted with two top-performing methods from Section 3.3: FrozenBiLM and LLoVi. Further details on the experiment can be found in Appendix E.
Our results, shown in Figure 7, reveal the following trends: (1) Both Language-only and Vision-Language settings show substantial improvements with larger temporal windows. For instance, in the language-only setting, LLoVi’s accuracy increases from 38.5% at the shot-level to 64.2% at the movie-level, reflecting a 25.7% gain. Similarly, FrozenBiLM shows an increasing trend but with a smaller gain in accuracy (+7.3%). The multimodal setting of LLoVi’s accuracy increases from 50.1% to 55.6% at the scene-level but shows only marginal gains at the movie-level. This plateau may be attributed to LLoVi’s naive fusion approach, which combines visual captions with subtitles, limiting further improvements; (2) The Vision-only modality exhibits low and stagnant performance across all levels suggesting that the current handling of visual data is inadequate for this task. Additional results for other methods are presented in Appendix E.
In conclusion, broader temporal windows significantly enhance SFD task performance, especially in language-only settings, underscoring the importance of long-form understanding in our design.
3.3 Baseline Comparison
In this section, we benchmark recent videoQA methods on SFD and evaluate 7 open-source models in three categories: FrozenBiLM [5], mPLUG-Owl2 [74] and Video-LLaVA [33] are VLMs that project visual frames into embedding of an LLM with the help of adapters; MovieChat [54] and TimeChat [47] have more advanced and specific settings, the former introduces a memory mechanism to handle long sequences and the latter proposes a timestamp-aware frame encoder for time-sensitive tasks; LLoVi [79] and LangRepo [26] are text-based methods: they both rely on visual captions on top of an LLM. LangRepo proposes an extra summarization and chain-of-thoughts [62] reasoning prompting.
These models are tested in a zero-shot video question-answering setting, adapted for long-form video understanding by incorporating subtitles and sampling the maximum number of frames possible. In the Multiple-Choice Question answering (MCQ) benchmark, we calculate the accuracy score based on the logits of the predicted options where applicable [5, 33, 74]. For models that generate plain-text answers, we evaluate them using a method similar to the Open-Ended Question answering (OEQ) where we rely on GPT-3.5 to compute the similarity between the predicted and correct answers, as detailed in [37]. All methods are evaluated across three modalities to better assess the contribution of each: Vision-Only (V) with video frames, Language-Only (L) with subtitles, and Vision-Language (VL) combining both. See Appendix D for more details. The results are summarized in Table 3.
User Study. To verify the answerability of our multiple-choice questions and assess the upper limit of SFD, we conducted three user studies accordingly to the aforementioned modality definitions: (1) Vision-Language (VL),—full video with audio and subtitles; (2) Vision-only (V)-muted videos; (3) Language-only (L)-plain text subtitles. For each question, all participants were asked to select the correct answer. Table 3 (last row) reports the results. More details on the user study are available in Appendix F. We observe that when provided with the full multimodal information, participants answer questions with high accuracy (89.8%). As expected, removing modalities lowers accuracy. Specifically, when using only subtitles the performance is 70.9%, whereas the vision-only performance drops to 59.0%.
Method | Venue | % Accuracy | |||||
Multiple-Choice QA | Open-Ended QA | ||||||
V | L | VL | V | L | VL | ||
Random | 20.0 | 20.0 | 20.0 | - | - | - | |
FrozenBiLM [5] | NeurIPS 2021 | 23.4 | 38.2 | 38.6 | - | - | - |
mPLUG-Owl2 [74] | CVPR 2024 | 38.3 | 20.7 | 21.3 | 22.1 | 1.8 | 1.6 |
Video-LLaVA [33] | arXiv 2023 | 34.2 | 21.3 | 24.7 | 19.2 | 6.4 | 8.0 |
LLoVi [79] | arXiv 2023 | 30.8 | 64.2 | 55.6 | 16.2 | 40.3 | 24.7 |
LangRepo [26] | arXiv 2024 | 29.0 | 32.1 | 31.0 | 3.5 | 10.4 | 9.5 |
MovieChat [54] | CVPR 2024 | 8.4 | 6.4 | 8.0 | 14.0 | 15.7 | 11.8 |
TimeChat [47] | CVPR 2024 | 25.5 | 6.4 | 31.8 | 26.4 | 9.4 | 5.9 |
Human | 59.0 | 70.9 | 89.8 | - | - | - |
Multiple-Choice Question answering (MCQ). These results highlight significant insights. Overall, model performance is generally poor, with LLoVi being a notable exception in the language-only setting (64.2%), driven largely by subtitles. Some models, such as FrozenBiLM, perform adequately in the vision-only setting (38.3%), but their performance drops when combining both modalities, indicating integration issues.
Compared to human performance, there is still substantial room for improvement. Accuracy is on par in the language-only setting, with only a 6.7% difference. However, as the best models mainly rely on subtitles, there is a significant gap in the multi-modal setting, where human performance reaches 89.8%. This discrepancy can be explained by the vision-only setting, where the gap reaches 20.7%. These results highlight that while models perform reasonably well with textual data, they struggle to effectively integrate visual information, underscoring the need for advancements in multi-modal fusion techniques to bridge the gap between human and model performance.
Open-Ended Question answering (OEQ). This task proves even more challenging, with lower accuracy across all models, ranging from 3.5% to 40.3%. Subtitles-only models, particularly LLoVi, lead in performance, highlighting the dominance of language processing in understanding and answering complex questions. These findings emphasize the difficulty of the task and the need for improved methods to handle and integrate multi-modal data effectively.
Overall, our results suggest that (1) text (in the form of subtitles) is a stronger cue than visual information for movie question-answering; (2) modern multimodal methods (max accuracy at 60.0%, see Table 3) fall behind human evaluation; and more importantly (3) there is a large room for improvement in the visual aspect, where the gap in performance between modern methods (average accuracy 38.3%) and user study (59.0% accuracy) is still very high.
4 Related Work
VideoQA benchmarks. Questions in such benchmarks can be framed to assess the reasoning, memory, and comprehension skills of both humans and algorithms. Hence, many datasets have been made available through the years, covering visual descriptions [69, 8], temporal action reasoning [65], compositional reasoning [19], social intelligence [76], instructional videos [71], egocentric videos [39, 24, 34], movies [56, 29, 38], among others [23, 67, 23, 32, 75, 63]. However, these datasets feature short videos and typically require reasoning on only a few frames to solve the task at hand, as shown by temporal certificates in [39].
VideoQA methods. To address video question answering, most modern approaches are based on large-scale pre-trained models, by deploying either multimodal contrastive learning [15, 31, 61, 78, 77, 10, 9] or visually-conditioned large language models [30, 11, 68, 74, 2, 70, 73]. For instance, mPLUG-Owl2 [68] enhances LLMs with visual capabilities, while FrozenBiLM [73], a frozen bidirectional language model, uses lightweight adapters and masked language modelling to excel in zero-shot tasks. FrozenBiLM excels in zero-shot video question-answering without using explicit supervision. More recently, Video-LLaVA [33] used a simple projection layer to connect a visual encoder and a large language model, fine-tuned with visual instruction tuning data.
Long-form video understanding. As new long-context models emerge [46], we need more complex benchmarks. Long-form video understanding evaluates the long-term reasoning capabilities of models. Some datasets have been introduced in recent years, mostly for egocentric videos [39] and movies [64, 53, 21, 56]. Instead, LVU [64] involves long movie clips but the subsequent tasks are relatively limited, focusing on irrelevant aspects like ’like’ ratio and view count prediction. [39] proposes multiple-choice questions on 3-minute videos that require on average 100 seconds of video watching to be answered correctly. [53, 21] introduce an audio description task, which is currently tackled at the clip level but could be extended over the whole movie. Finally, closer to our work, [56] and [54] propose a movie question-answering datasets focused on movie plots, settings and characters. However, they do not release any video data because of copyright issues.
5 Discussion
We introduced the Short Film Dataset (SFD), a long-form video understanding benchmark featuring complete and narrative-driven amateur short movies. It includes two types of QA tasks, multiple-choice and open-ended, that are automatically generated by LLMs and manually curated by human annotators. Compared to other long-form video datasets, SFD stands out by offering richer, story-oriented content with tasks specifically tailored for long-term reasoning, as validated by our experiments. Unlike most movie datasets that use copyrighted commercial films prone to data leakage with LLMs, SFD relies on amateur films, which are publicly accessible and have a limited online presence. Furthermore, our analysis indicates that while language-based methods achieve performance levels comparable to humans, state-of-the-art vision-based and multimodal methods fall behind human evaluation and they still have considerable room for improvement, highlighting the need for further advancements. In conclusion, we believe that our SFD paves the way for accessible, and comprehensive long-term movie understanding that is not known a priori by LLMs, thus helping the community develop robust methods. Our dataset and code will be made publicly available.
6 Acknowledgement
This work was supported by ANR-22-CE23-0007, Hi!Paris grant and fellowship, and was granted access to the High-Performance Computing (HPC) resources of IDRIS under the allocations 2023-AD011014489 made by GENCI. We would like to thank Nicolas Dufour, Robin Courant, Pierre Vassal, and the anonymous reviewers for their insightful comments and suggestions. We also express our gratitude to the annotators of the human study for their help.
Appendix
In this appendix, we show additional materials of the paper Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding. We present the following items:
Appendix A Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes]
Did you discuss any potential negative societal impacts of your work? [Yes] See Supplementary Materials Section B
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Supplementary Materials Section B
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://huggingface.co/datasets/rghermi/sfd
- (b)
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Our experiments consist of benchmarking various aspects of our dataset. For a given compute, instead of performing one experiment several times (with different random seeds), we chose to experiment with several different multimodal methods and multiple LLMs. We observed that they all report similar performances and overall follow the same trend; hence, we argue that our findings are generalizable.
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 3.3
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] See Section B.
Did you mention the license of the assets? [Yes] See Section B.
Did you include any new assets either in the supplemental material or as a URL? [N/A]
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section B.
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section F.
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix B Ethical considerations
1.
Copyright and License:
We do not distribute raw video content; instead, we provide only URLs redirecting to YouTube, where the full copyright and licensing rights of creators are acknowledged. Additionally, by sharing the URLs, YouTube’s Terms of Service regarding exclusive rights are respected.333https://www.youtube.com/static?template=terms (Section: License to Other Users).
The dataset will be used solely for academic research purposes. To further protect the metadata and other information, we apply the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license to our dataset.
2.
Potential negative social impact:
Bias: Collected movies are primarily from North America and Europe, mainly English-speaking, the under-representation of the content towards non-English-speaking population and culture, potentially leading to unbalanced, biased video understanding model training and testing.
Impact on creative process: Progress on movie understanding may encourage filmmakers to tailor their work to optimize for what is favoured by algorithms, potentially stifling creative freedom and depriving innovation.
Appendix C Data Preprocessing and Statistics
SFD | |
#Movies | 1,078 |
#Clips | 161,321 |
#Clips / Movie | 151 |
Avg. Clip Len. (s) | 5.43 |
#Captions | 161,321 |
#Location Tags | 161,321 |
SFD | |
#Face Instances | 56,265 |
#Unique Faces | 2,598 |
#Unique Faces / Movie | 2.41 |
#Subtitles | 23,820 |
SFD, alongside the videos and subtitles (see Section 2), also offers a comprehensive set of annotations: clips (shots), shot captions and scales, location tags, and face tracks. These resources, together with tools for downloading, preprocessing, and annotation, will be made available as part of the dataset release. Below, we detail their creation processes.
Clips (Shots). In the context of movies, a clip corresponds to a shot, characterized by its composition, length, and camera movement. We use PySceneDetect444https://www.scenedetect.com/ to segment movies in SFD into 161,321 shots, with each movie containing an average of 151 shots, as shown in Table 5(a). The number of shots varies by genre; for example, comedies average 163 shots, while dramas have about 146.
Clip information. For each shot, we extract detailed shot-level information. We generate shot captions using image and video captioning tools (LLaVA [36], BLIP-2 [30], Video-LLaVA [33]), and we identify locations using CLIP features [43] based on a curated list of locations from MovieCLIP [6].
Face Tracks. We use an advanced facial recognition pipeline 555https://github.com/serengil/deepface to identify over 56k+ face instances in the dataset, corresponding to 2.6k+ unique characters, with an average of 2.41 characters per movie, as shown in Table 5(b).
These annotations enrich the dataset, providing deeper insights for tasks related to character analysis and interaction within the films.
In Figure 8, we perform a word cloud analysis for three video domains commonly used in the literature: instructional videos [71], egocentric videos [39], and movies (ours). We notice that instructional and egocentric videos are focused on actions (verbs like make, go, cut, use, pick, put) and objects (e.g. table, sugar, oil, thing, stuff, table, paper), with an emphasis on body parts for egocentric videos (right hand, left hand). For movies, the vocabulary is more focused on characters (e.g. man, woman, boy, girl, young, old), relationships (e.g. mother, father, family, couple), time (e.g. day, night, year), and locations (e.g. home). These clearly show the story component of SFD.
Appendix D Technical details on baselines
In this section, we present additional details and results for the baselines discussed in Section 3.3 and Table 3 main paper, where we report the benchmark performance of the state-of-the-art video understanding methods on our SFD.
FrozenBiLM [5]. We use FrozenBiLM pre-trained on WebVid10M[5]. Following[39], we reformulate the question-answering task for this model with the following prompt: ‘Question: {question} Is it {option_i}?’. We get logits for the yes/no tokens for each option and take the maximum as the final prediction. We report results with and without subtitles. We compute the accuracy score, by comparing the predicted option index and the correct option index. As it is a masked language model, FrozenBiLM is not suited to open-ended question answering.
mPLUG-Owl2 [74]. We use the instruction-tuned version (‘MAGAer13/mplug-owl2-llama2-7b’), based on ViT-L and LLaMa-7b. Inference is the same as FrozenBiLM. In the open-ended question-answering setting, we only provide the question and record a free-form text output: ‘Question: {question} Please answer with a short sentence.’. Following [37], we use GPT-3.5 (‘gpt-3.5-turbo’) to compute the similarity between the predicted and the correct answer. We report this LLM-based accuracy.
Video-LLaVA [33]. We use the fine-tuned version (‘LanguageBind/Video-LLaVA-7B-hf’), based on LanguageBind and Vicuna-13B v1.5. Inference is the same as mPLUG-Owl2, in both multiple-choice and open-ended question-answering settings.
LLoVi [79]. Following the official implementation, we use LLaVA (‘llava-hf/llava-1.5-13b-hf’) to extract captions at 0.5 FPS and GPT-3.5 (‘gpt-3.5-turbo’) as a backbone LLM. We experiment with the prompting strategy proposed in the original paper [79]. The prompt for multiple-choice question answering is as follows:
‘Please provide a single-letter answer (A, B, C, D, E) to the following multiple-choice question, and your answer must be one of the letters (A, B, C, D, or E). You must not provide any other response or explanation. You are given some language descriptions of a video. The video is {duration} seconds long. Each sentence describes a {clip_length}s clip. The descriptions are sequential and non-overlapping which cover the whole video exactly. Here are the descriptions: {captions}. You are going to answer a multiple-choice question based on the descriptions, and your answer should be a single letter chosen from the choices. Here is the question: {question}. Here are the choices. A: {option_0} B: {option_1} C: {option_2} D: {option_3} E: {option_4}")’
LangRepo [26]. Similar to LLoVi, LangRepo is another LLM-backed video understanding method. We use the open-sourced official implementation, Mixtral 87B (‘mistralai/Mixtral-8x7B-Instruct-v0.1’). Unlike LLoVi, LangRepo operates in two stages: The first is summarization, where LangRepo consolidates similar information from descriptions, captions, subtitles, or both, using specific prompts following the official implementation:
‘You are a helpful expert in video analysis. You are given a list of {num_to_rephrase} language descriptions for a video. Each sentence describes a {clip_length}s clip. Here are the descriptions as a list:{memory}. Please summarize and rephrase each item in the list as a single sentence of {num_words_in_rephrase} words. Keep the same original subject. Keep all information intact without leaving anything out. Return only the rephrased list of {num_to_rephrase} descriptions in the same order, without additional details.’
In the second stage, the question-answering is based on the rephrased summarization with prompts:
‘Here are the descriptions: {narration}. You are going to answer a multiple-choice question based on the descriptions. Here is the question: {question}? You must select one of these choices as the answer: A: {option_0} B: {option_1} C: {option_2} D: {option_3} E: {option_4} The correct answer is:’
MovieChat [54]. We use the official implementation of MovieChat (‘rese1f/MovieChat’, which incorporates LLaMa2-7B (‘meta-llama/Llama-2-7b’ and corresponding Video-LLaMa-7B (‘DAMO-NLP-SG/Video-LLaMA-2-7B-Finetuned’). Similarly, we use GPT-3.5 (‘gpt-3.5-turbo’) [42] to compute the accuracy of responses. Following the official implementation, we use the following prompt:
‘By watching the video and combining it with the caption/subtitle from video , you are required to answer a question with multiple options: Question: {question}; Option A: {option_0}; Option B: {option_1}; Option C: {option_2}; Option D: {option_3}; Option E: {option_4}; Answer to the multiple-choice question based on the textual information extracted from the movie. If you don’t know the answer, please output ‘Unknown’. Your output should be just one of A, B, C, D, E and nothing else.’
TimeChat [47]. We use the official implementation of TimeChat (‘ShuhuaiRen/TimeChat-7b’). Following the official implementation, the prompt and the accuracy computation details are the same as those in the previous MovieChat model.
For all models, we also incorporate subtitles as input. For that, we concatenate each query with the following text: ‘Subtitles: {subtitles}’. Subtitles are formatted as follows: ‘{start_timestamp} - {end_timestamp} {subtitle_i}’. We trim subtitles if they cannot fit into the input size. Results and analysis can be seen in Section 3.3 and Table 3.
Appendix E Technical details on the temporal window study
In this section, we present additional details and results for the temporal window study (see Section 3.2 and Figure 7). The goal of the temporal window study is to evaluate the performance of various methods on SFD at different temporal levels—shot, scene, and movie—across all modalities. Along with the discussed FrozenBiLM and LLoVi, we examine Video-LLaVA and LangRepo.
At the shot-level, we truncate the input prompt to include only information from a specific shot. For FrozenBiLM and Video-LLaVA, it means that we only include subtitles and frames extracted from the shot. For LLoVi and LangRepo, we only include subtitles and captions extracted from the shot. Hence, we get logits for each answer option and shot, aggregated in a matrix of shape : we obtain a final prediction by taking the maximum logit over both dimensions. Note that the shot-level inference takes too much time to compute for LangRepo ( A100 GPU-hours per experiment), as this method involves many steps of rephrasing and summarizing.
The scene-level inference is similar to the shot-level one, the main difference being that we constrain the input information (subtitles and frames/captions) to the scene-level, which arbitrarily corresponds to 10 shots in our case. We end up with a logit matrix of size .
The movie-level inference is straightforward: it is the same as in Section D.
Vision | Language | Vision-Language | ||
FrozenBiLM | Shot-Level | 22.1 | 30.9 | 30.8 |
Scene-Level | 22.7 | 34.8 | 34.8 | |
Movie-Level | 23.4 | 38.2 | 38.6 | |
Video-LLaVA | Shot-Level | 16.0 | 24.4 | 24.0 |
Scene-Level | 23.7 | 20.1 | 20.1 | |
Movie-Level | 34.2 | 21.3 | 24.7 | |
LLoVi | Shot-Level | 32.8 | 38.5 | 50.1 |
Scene-Level | 34.2 | 51.2 | 55.4 | |
Movie-Level | 30.8 | 64.2 | 55.6 | |
LangRepo | Shot-Level | - | - | - |
Scene-Level | 26.9 | 27.9 | 28.1 | |
Movie-Level | 29.0 | 32.1 | 31.0 | |
Human | Movie-Level | 59.0 | 70.9 | 89.8 |
The numerical results of the temporal window study are reported in Table 6.
Along the temporal direction (i.e. row-wise for each method), we observe that: (i) generally, all methods exhibit consistent gains with larger temporal windows confirming the importance of long-term understanding for our designed questions, and (ii) exceptions are noted, such as with Video-LLaVa in Language and Vision-Language settings, likely due to its low precision (near random chance at 20%) which diminishes the impact of the temporal window.
For different modalities (i.e. column-wise), we make the following observations: (i) Corroborating the argument in the main paper (Section 3.3), the Language-only models, in general, achieve better performances, e.g. LLoVi reaches the best performance at movie-level in the language-only setting (64.2%). (ii) The vision-only setting results in the weakest performance, revealing the largest gap compared to human performance (last row). This suggests significant potential for improvement in visual models. (iii) Current modality-fusion methods show limited effectiveness. VLM models like FrozenBiLM and Video-LLaVA only demonstrate marginal improvements when fusing vision and language, with performance increment from 38.2% to 38.6% for FrozenBiLM and from 21.3% to 24.7% for Video-LLaVa. However, all LLM-based methods experience a regression after modality fusion.
Overall, the human benchmark highlights that there remains a substantial gap between the best methods and human performance, particularly in vision-language integration over extended temporal windows.
Appendix F Human Study
In this section, we present additional details for the conducted human study (see Section 3.3, ‘User Study’). We conducted a comprehensive human study to evaluate the upper-limit performance of our benchmark. Participants were asked to watch a movie and subsequently answer multiple-choice questions based on the content. To analyze the impact of different modalities, we implemented two ablation conditions: a muted version, where participants had access to the vision-only information without audio or subtitles, and a blind version, where participants were provided with the audio-only information, in the form of textual subtitles, without any video.
To avoid any bias, we ensured that different sets of participants were assigned to the full, vision-only, and audio-only settings, preventing any overlap and potential influence from experiencing multiple conditions. Movies were randomly assigned to users to ensure a diverse range of data. The entire study was facilitated through an integrated interface, which restricted access to external information such as movie loglines and synopses, with only the movie title being displayed (see Figure 9).
Appendix G Technical details on data leakage experiment
Name | SFD | MovieQA | LVU |
Open-source models | |||
Gemma 2B | 19.7 | 18.3 | 28.8 |
Gemma 7B | 20.7 | 21.5 | 52.4 |
LLaMA 3 8B | 22.1 | 34.5 | 69.9 |
LLaMA 3 70B | 24.1 | 51.9 | 64.1 |
Mistral 7B | 28.9 | 44.1 | 61.4 |
Mixtral 8x7B | 33.5 | 55.4 | 70.2 |
Commercial models | |||
Claude 3 Haiku | 28.9 | 55.4 | 68.5 |
Claude 3 Sonnet | 36.0 | 64.4 | 71.0 |
GPT-3.5 | 26.3 | 56.7 | 75.0 |
GPT-4 | 31.5 | 71.3 | 76.0 |
In this section, we include additional information on the data leakage study mentioned in the main paper Section 3.1. The objective of our data leakage study is to demonstrate that current large language models (LLMs) possess prior knowledge about commercial movies. This is because LLMs have been exposed to vast amounts of movie-related textual information during pretraining, including synopses, blog posts, news articles, and more. The experiment is set up as an open-ended question-answering task where the input consists of a movie title and a corresponding question, and the output is a free-form text generated by the LLM. We then evaluate the answer using an LLM to compute the similarity between the predicted and the correct answer, following [37].
We choose ten of the most recent LLMs, among which six are open-source models: Gemma 2B, Gemma 7B, LLaMA-3 8B, LLaMA-3 70B, Mistral 7B, Mixtral 8x7B; and four commercial models: Claude 3 Haiku, Claude 3 Sonnet, GPT-3.5, GPT-4. The open-source models are run using the Hugging Face transformers library 666https://github.com/huggingface/transformers. The commercial models are run using their official API: the OpenAI API 777https://openai.com/api for GPT and the Anthropic API 888https://www.anthropic.com/api for Claude 3.
We choose three movie datasets: MovieQA [56], LVU [64], and SFD (ours). For MovieQA and SFD, the question-answer pairs are directly available. In the case of LVU, we re-formulated the classification tasks into templated questions to fit the open-ended question-answering format. For example, the director classification task is converted into the question ‘Who directed the movie {movie_title}?’.
The detailed results of the data leakage experiment are presented in Table 7 and Figure 10. Their analysis shows that LLMs exhibit significantly higher data leakage on MovieQA and LVU datasets compared to SFD. Models like GPT-4 and Claude 3 Sonnet answer correctly to more than 70% of the questions of both datasets, without any context except the movie title. The scatter plot emphasizes a correlation between data leakage and model knowledge (i.e. MMLU performance), with higher MMLU scores correlating with increased leakage issues on compromised datasets. This trend is consistent across both open-source and commercial models, highlighting the vulnerability of larger and more advanced models to memorizing training data.
In sharp contrast, the relatively low and flat performance trend on SFD suggests it is a more robust benchmark for evaluating model capabilities. These findings emphasize the need for robust evaluation datasets to accurately assess model performance and mitigate the impact of data leakage.
Appendix H Samples from SFD
In the following pages, we show some samples from SFD. For each sample, we display the title, the logline, the synopsis, a few frames, some subtitles, and the designed multiple-choice questions. See Figures 11, 12,13, 14.
Appendix I Prompts
We provide the prompts for the data generation process: question generation prompt (see Figure 8), distractor generation prompt (see Figure 9), and question classification prompt (see Figure 10). As detailed in Section 2.2 , those prompts were used with GPT-4 (’gpt-4-turbo-2024-04-09’) through the OpenAI API.
![[Uncaptioned image]](x20.png)
![[Uncaptioned image]](x21.png)
![[Uncaptioned image]](x22.png)
Appendix J Datasheet for Datasets
Motivation
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
This dataset and benchmark are created to assess long-form video understanding capabilities of modern multimodal systems. It is different from other video question-answering datasets as it features longer videos and associated story-related questions. Additionally, it stands out as a dataset that is available online and not exposed to current large language models (limited data leakage issues).
Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
This dataset was created by Ridouane Ghermi††footnotemark: , Xi Wang††footnotemark: , Vicky Kalogeiton††footnotemark: , and Ivan Laptev999MBZUAI.
Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
This project is funded by V.Kalogeiton ANR-22-CE23-0007 and a Hi!Paris project.
Any other comments?
Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
Our dataset instance consists of a short movie (around 5-20 mins), some metadata (e.g. movie title, logline, synopsis, country of origin, language, release year), and a list of multiple-choice questions and answers.
How many instances are there in total (of each type, if appropriate)?
There are 1,078 movies and 4,885 questions. Each question comprises one correct answer and four wrong answers.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
The dataset is not a sample of instances from a larger set.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
Each instance includes a movie as a video file. Metadata comes as .csv files, in the form of text. We also provide processed features.
Is there a label or target associated with each instance? If so, please provide a description.
Each movie comes along with several multiple-choice questions, each with five possible options. The correct option is designated by an index (from 0 to 4).
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g. because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
No.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
We provide a dataframe that links question ids and video ids.
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
SFD is a test set only.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
No, the dataset has been manually curated.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
All movies are available on YouTube; We only provide the corresponding URLs. It should remain constant over time as it comes from YouTube channels that curate a movie catalogue. The dataset is distributed under the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals non-public communications)? If so, please provide a description.
No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
Yes, the dataset contains movies with actors.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
No.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
Yes, it is possible to identify actors.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
No, everything is fictional.
Any other comments?
Collection Process
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
Movies are available online and downloaded from YouTube using yt-dlp, along with metadata. Multiple-choice questions are generated by a commercial LLM, GPT-4, and further manually curated.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
Video URLs are scraped using the official YouTube API. Videos and metadata are downloaded using yt-dlp. Questions are generated using GPT-4 through the official OpenAI API and further refined by manual curation.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
Not applicable.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
Students voluntarily helped to curate the dataset.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
All data are collected and curated during the first half of 2024. Movies were uploaded to the video-sharing platform between 2017 and 2024.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
No.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
No.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
Not applicable.
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
Not applicable.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Not applicable.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
Not applicable.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
Not applicable.
Any other comments?
Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
We extracted several annotations from the dataset, see Sections C.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
Yes.
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
Yes. For annotation extraction, we will provide the tools and scripts in the GitHub repository. For question generation, we will provide the prompts but the usage requires an access to the OpenAI API (which is not free).
Any other comments?
Uses
Has the dataset been used for any tasks already? If so, please provide a description.
The dataset is accompanied by two tasks: multiple-choice and open-ended question answering.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
Yes, we will make this available on the dataset GitHub repository.
What (other) tasks could the dataset be used for?
We strongly believe that, with appropriate annotations, this dataset can be extended for many challenging use cases, e.g. spatio-temporal localization, video grounding, movie summarization, causal reasoning.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
No.
Are there tasks for which the dataset should not be used? If so, please provide a description.
No.
Any other comments? –
Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
Yes, but we do not distribute raw video content; instead, we provide only URLs redirecting to YouTube, where the full copyright and licensing rights of creators are acknowledged and YouTube’s Terms of Service regarding exclusive rights are respected. To further protect the metadata and other information, we apply the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license to our dataset.
How will the dataset be distributed (e.g., tarball on website, API, GitHub) Does the dataset have a digital object identifier (DOI)?
The dataset is distributed on the HuggingFace platform and as a GitHub repository. We provide video URLs and annotations. Additionally, we provide code to download videos, pre-process the data and run baselines.
When will the dataset be distributed?
The dataset is already available.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
The dataset is distributed under the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license. The data was collected from publicly available sources. All rights and credits for the original movie content go to the respective owners. This dataset and any derivatives are intended for non-commercial use only and must be shared under the same license.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
Not applicable.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
Not applicable.
Any other comments?
Maintenance
Who will be supporting/hosting/maintaining the dataset?
The dataset is hosted on both GitHub and HuggingFace. It will be maintained by the authors.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
We provide a contact information email, and will also answer any pull requests in the repository.
Is there an erratum? If so, please provide a link or other access point.
The erratum is available on the GitHub repository.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
The dataset is a test benchmark, it will remain stable to ensure fair evaluation of future methods.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
Not applicable.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Not applicable.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
If others want to contribute to the dataset, by adding more tasks for example, we would gladly host them in our repository. Also, we will provide as many details and tools as possible to reproduce our annotation process and create more tasks.
Any other comments?
N/A.

### v2

@@
-Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding
+Long Story Short: Story-level Video Understanding from 20K Short Films
 Abstract
-Recent advances in vision-language models have significantly propelled video understanding.
+Recent developments in vision-language models have significantly advanced video understanding.
 Existing datasets and tasks, however, have notable limitations.
 Most datasets are confined to short videos with limited events and narrow narratives.
-For example, datasets with instructional and egocentric videos often document the activities of one person in a single scene.
-Although some movie datasets offer richer content, they are often limited to short-term tasks, lack publicly available videos and frequently encounter data leakage given the use of movie forums and other resources in LLM training.
-To address the above limitations, we propose the Short Film Dataset (SFD) with 1,078 publicly available amateur movies, a wide variety of genres and minimal data leakage issues.
-SFD offers long-term story-oriented video tasks in the form of multiple-choice and open-ended question answering.
-Our extensive experiments emphasize the need for long-term reasoning to solve SFD tasks.
-Notably, we find strong signals in movie transcripts leading to the on-par performance of people and LLMs.
-We also show significantly lower performance of current models compared to people when using vision data alone.
+For example, datasets with instructional and egocentric videos often depict activities of one person in a single scene.
+Although existing movie datasets offer richer content, they are often limited to short-term tasks, lack publicly available videos, and frequently encounter data leakage issues given the use of subtitles and other information about commercial movies during LLM pretraining.
+To address the above limitations, we propose Short-Films 20K (SF20K), the largest publicly available movie dataset.
+SF20K is composed of 20,143 amateur films and offers long-term video tasks in the form of multiple-choice and open-ended question answering.
+Our extensive analysis of SF20K reveals minimal data leakage, emphasizes the need for long-term reasoning, and demonstrates the strong performance of recent VLMs.
+Finally, we show that instruction tuning on the SF20K-Train set substantially improves model performance, paving the way for future progress in long-term video understanding.
 1 Introduction
@@
 Despite their remarkable progress, current datasets are still constrained by some limitations.
-Most of them consist of short videos [20, 1, 13, 69, 72, 17, 55, 27, 32, 19, 65], often under one minute, and primarily target short-term tasks e.g., action recognition and video retrieval, often requiring only a few seconds of visual content for problem-solving [39, 28].
-For long-form video understanding, the community has turned to three video categories.
-(i) Egocentric videos, an emerging domain featuring extended first-person sequences that capture a person’s continuous actions [18, 12, 24, 34, 39, 52, 51] or chronicle daily activities, offering tasks from object manipulation to social interactions [18].
-Nevertheless, while such datasets excel in video duration, they lack narrative depth, focusing mainly on immediate visuals rather than rich storytelling; (ii) Instructional videos [41, 71], which depict a wide variety of human tasks and focus on the procedural aspect, are often brief and deficient in storytelling; (iii) Movies, unlike egocentric and instructional videos, provide complex plots and extended duration, making them an ideal testing ground for developing models that can comprehend long-form video narratives, see Figure 1.
-Several movie datasets have been proposed in the literature [48, 4, 56, 29, 53, 21, 64, 22, 59, 66, 49, 40, 44, 7, 6, 38] offering rich narrative-driven content and various tasks.
-These tasks include clip retrieval [4], video question answering [56, 29, 38], audio description [53, 21], semantic role labeling [49], scene segmentation [22, 44], and visual scene recognition [6].
-However, current movie datasets suffer from three main limitations.
-(i) Accessibility: these datasets are not publicly available as they typically contain commercial copyrighted movies (see Figure 2).
-(ii) Clip duration: they often consist of incomplete and short clips [56, 29, 45, 64] (i.e., less than 4 minutes, see Figure 2), limiting their potential for long-form understanding with evolving storylines.
-Moreover, these datasets frequently suffer from ill-designed benchmark questions with insufficient narrative focus [54].
-(iii) Data Leakage: as these datasets comprise well-known commercial movies, modern Large Language Models (LLMs) and Vision-Language Models (VLMs) have likely been exposed to some form of movie information (i.e., synopses, reviews, discussions, subtitles, blog posts).
-Figure 3 showcases this, where given only movie titles, modern LLMs can obtain high accuracy when answering questions from main movie datasets.
-This data leakage leads to biased benchmarks and ineffective training.
-To address the above limitations, we introduce the Short Film Dataset (SFD), a novel video benchmark comprising 1,078 short films, totalling over 243 hours of videos with an average duration of 13 minutes per film—significantly longer than existing datasets (see Figure 2).
-Unlike copyrighted films, SFD includes publicly accessible amateur films of diverse genres from YouTube.
-Although these films are shorter than traditional movies, they feature complex narratives that unfold through key sequences of events and multiple character interactions.
-Most importantly, with minimal exposure to LLMs, our dataset is less prone to data leakage issues plaguing other movie datasets (see Figure 3).
-Aiming towards long-term story-level multimodal video understanding, we propose two question-answering tasks on SFD: (a) following recent trend [56, 45], we propose Multiple-Choice Question answering (MCQ); (b) We also introduce Open-Ended Question answering (OEQ), a more generative and unrestricted QA format that more realistically and objectively measures a model’s ability to comprehend long-form videos.
-All questions and answers are generated by LLMs from movie descriptions and undergo thorough manual curation, ensuring they accurately reflect the settings, characters, storylines, and themes of the movies.
-To validate the advantages of SFD, we conduct extensive experiments and quantify data leakage, the benefits of long-term reasoning and comparison to human performance while using state-of-the-art vision and language models.
-Our findings are: (i) Unlike existing VideoQA datasets, SFD provides story-level tasks defined on public data with minimal data leakage (Section 3.1); (ii) our long-term reasoning experiments in Section 3.2 confirm the advantage of movie-level video understanding for solving SFD tasks; (iii) the evaluation in Section 3 reveals a large gap between recent VideoQA methods and human performance in vision-only and multimodal settings.
-SFD, hence, provides a unique benchmark for evaluating and advancing new methods for story-level long-term video understanding.
-Our dataset, code, and models are publicly available at https://shortfilmdataset.github.io.
-2 Short Film Dataset
-Short films are motion pictures that typically last between 5 and 20 minutes.
-As illustrated in Figure 4, short films span a variety of styles (e.g., narrative fiction, documentary, animation, etc.) and genres (e.g., action, drama, comedy, horror, etc.).
+Most of them consist of short videos [23, 1, 15, 76, 78, 20, 60, 29, 34, 22, 71], typically under one minute, and primarily target short-term tasks such as action recognition and video retrieval, which often require only a few seconds of visual content to solve [41, 30].
+Long-form video understanding, on the other hand, has been mainly studied in three domains: (i) Egocentric videos, featuring extended first-person sequences with continuous actions [21, 13, 27, 37, 41, 57, 56] and daily activities ranging from object manipulation to social interactions [21].
+Although these datasets excel in video duration, they lack narrative depth, focusing on immediate visuals rather than rich storytelling; (ii) Instructional videos [44, 77], which depict a variety of human tasks and focus on the procedural aspect, are often brief and deficient in storytelling; (iii) Movies, unlike egocentric and instructional videos, provide complex plots and extended durations, making them well suited for developing models capable of comprehending long-form narratives (see Figure 1).
+Several movie datasets have been proposed in the literature [53, 5, 61, 31, 58, 24, 70, 25, 64, 73, 54, 42, 48, 8, 7, 40, 59, 50, 72] offering rich, narrative-driven content and supporting various tasks.
+These tasks include clip retrieval [5], video question answering [61, 31, 40, 59, 50], audio description [58, 24, 72], semantic role labeling [54], scene segmentation [25, 48], and visual scene recognition [7].
+However, current movie datasets suffer from three main limitations: (i) Accessibility: access to commercial movies is subject to copyright restrictions (see Figure 2); (ii) Clip duration: they often consist of incomplete and short clips [61, 31, 49, 70], often less than four minutes in length (see Figure 2), limiting their potential for long-form understanding of evolving storylines.
+Moreover, these datasets frequently suffer from ill-designed benchmark questions with insufficient narrative focus [59]; (iii) Data leakage: as these datasets feature well-known commercial movies, modern Large Language Models (LLMs) and Vision-Language Models (VLMs) have likely been exposed to related content, such as synopses, reviews, discussions, subtitles, and blog posts.
+Figure 3 showcases this issue, showing that modern LLMs can achieve high accuracy on questions from major movie datasets using only movie titles.
+This leakage leads to biased benchmarks and ineffective training.
+To address the limitations of existing movie datasets, we introduce Short-Films 20K (SF20K), a novel video dataset comprising 20,143 short films totaling almost 3,582 hours of video, with an average duration of 11 minutes per film — significantly longer than existing datasets (see Figure 2).
+Unlike copyrighted films, SF20K consists of publicly accessible amateur films across diverse genres sourced from YouTube and Vimeo.
+While shorter than traditional movies, these films feature complex narratives that unfold through key sequences of events and multiple character interactions.
+Crucially, with minimal exposure to LLMs, SF20K is less prone to data leakage issues plaguing other movie datasets (see Figure 3).
+Aiming towards long-term story-level video understanding, we propose two question-answering tasks on SF20K: (a) Multiple-Choice Question Answering (MCQA), following recent trends [61, 49]; and (b) Open-Ended Question Answering (OEQA), a less restrictive QA format requiring free-form answers.
+The questions and answers are initially generated using LLMs based on movie descriptions.
+For the test set, all QAs undergo meticulous manual curation to ensure they accurately reflect the settings, characters, storylines, and themes of the films.
+To demonstrate the benefits of SF20K, we conduct extensive experiments including an analysis of data leakage, the impact of long context windows, model performance relative to human benchmarks using state-of-the-art vision-language models, and the effect of instruction tuning on the training set.
+Our contributions are threefold: (1) We propose SF20K, the largest publicly accessible movie dataset, comprising 20,143 short films and offering two story-level videoQA tasks.
+(2) Through extensive experiments, we show: (i) Unlike other movie datasets, SF20K exhibits minimal data leakage to modern LLMs; (ii) Recent video models show significantly lower performance compared to humans, particularly in the multimodal setting; (iii) A long temporal window is required to solve the QA tasks, highlighting the need for models capable of handling information at the movie level.
+(3) We demonstrate that large-scale instruction tuning on SF20K-Train helps model performance.
+We argue that SF20K provides a unique dataset for advancing long-term, story-level video understanding.
+The dataset, code, and models are publicly available at https://ridouaneg.github.io/sf20k.html.
+2 Related Work
+Video QA benchmarks.
+Questions in such benchmarks can be framed to assess the reasoning, memory, and comprehension skills of both humans and models.
+Hence, many datasets have been made available through the years, covering visual descriptions [76, 9], temporal action reasoning [71], compositional reasoning [22], social intelligence [82], instructional videos [77], egocentric videos [41, 27, 37], movies [61, 31, 40], among others [26, 74, 26, 34, 81, 69].
+However, these datasets feature short videos and typically require reasoning on only a few frames to solve the task at hand, as shown by temporal certificates in [41].
+Long-form video understanding.
+As new long-context models emerge [51], we need more complex benchmarks.
+Long-form video understanding evaluates long-term reasoning capabilities of those models.
+Some datasets have been introduced in recent years, for egocentric videos [41], movies [70, 58, 24, 61, 50], and general videos [17, 66].
+For instance, [41] proposes multiple-choice questions on 3-minute videos that require on average 100 seconds of viewing to be answered correctly.
+[70] includes long movie clips, however the subsequent tasks are relatively limited, focusing on less relevant aspects such as ’like’ ratio and view count prediction.
+[58, 24] introduce an audio description task, which is currently tackled at the clip level but could be extended over the whole movie.
+Recently, [17, 66] introduced longer videos that present significant challenges even for the most advanced vision-language models.
+[61, 59] propose movie question-answering datasets focused on movie plots, settings and characters, though access to the data is limited.
+Finally, closer to our work, [50] offers a dataset of question-answer pairs derived from commercial movie clips, providing both training and test sets; however, along with other movie datasets, it may suffer from data leakage issues.
+3 SF20K: A Dataset of 20K Short Films
+Short films are motion pictures that typically last between 5 and 40 minutes.
+As illustrated in Figure 4, short films span a variety of styles (e.g., narrative fiction, documentary, animation) and genres (e.g., drama, comedy, horror).
 Filmmakers often use the short film format to explore new ideas, techniques, and storytelling styles.
 Short films, hence, have more of an experimental rather than commercial value and are often made publicly available.
-Samples from the dataset can be found in Appendix H.
-The amount of publicly available short films has significantly increased over the last years (see Figure 6(c)).
-We take advantage of this development and create a Short Film Dataset (SFD) to foster research on story-level video understanding.
-In the remaining part of this section, we describe the procedure for collecting videos and corresponding metadata (Section 2.1) as well as generating question-answer pairs (Section 2.2).
-Section 2.3 presents the analysis and statistics of SFD.
-The overall pipeline for SFD creation is illustrated in Figure 5.
+The amount of publicly available short films has significantly increased over the last few years (see Figure 5(b)).
+We take advantage of this development and create SF20K, a dataset of 20,143 short movies collected from online platforms, totaling 3,582 hours of video content, with each video lasting on average 11 minutes.
+Our procedure for collecting videos, metadata, and annotations for SF20K is described in Section 3.1.
+To promote the development of long-form, story-oriented video understanding, we introduce two video QA tasks: Multiple-Choice Question Answering (MCQA) and Open-Ended Question Answering (OEQA).
+The QA generation process — consisting of synopsis preparation, automatic question generation, and manual curation — is presented in Section 3.2.
+Section 3.3 presents detailed statistics about SF20K, including annotations, tasks, and comparison to other datasets.
+Section 3.4 presents a specific analysis of SF20K-Test, including question types and subsets.
 3.1 Data Collection
-To build our dataset, we capitalize on the abundance of short films on video-sharing platforms.
-We specifically target the Omeleto YouTube channel111https://www.youtube.com/@Omeleto and its subchannels, including @OmeletoComedy and @OmeletoDrama, among others.
-These channels promote high-quality and award-winning short films, some of which have been recognized by Oscars and BAFTA awards.
-We download videos, subtitles, and associated metadata from YouTube using yt-dlp222https://github.com/yt-dlp/yt-dlp.
-As speech transcripts were originally available only for a portion of the dataset, we have used WhisperX [3] to extract missing transcripts.
-The metadata accompanying each short film includes the movie title, a logline which is a concise one-sentence summary of the plot, and additional movie details such as genre, release year, region of origin, and language.
-Each film also comes with a detailed description that offers a deeper understanding of various elements such as a synopsis providing a brief narrative overview, the director’s inspirations, storytelling approach, actors’ performances, and other external information.
-To maintain clarity and objectivity in our dataset, we retain only the synopsis and exclude other subjective or ambiguous elements of the description.
-2.2 Question and Answer Generation
-To promote the development of long-form and story-oriented video understanding, we complement SFD with two Video Question Answering (VideoQA) tasks: Multiple-Choice Question answering (MCQ) and Open-Ended Question answering (OEQ).
-We first automatically generate question-answer pairs by prompting GPT-4 [42] with movie titles, loglines and synopses.
-We then manually curate questions and answers to ensure their relevance and correctness.
-We describe the generation process in more details below.
-Automatic QA pair generation.
-For each film we obtain human-written information, i.e.
-movie titles, loglines and synopses from Omeleto YouTube channels.
-We then prompt GPT-4 [42] to propose relevant question-answer pairs from movie information, with dedicated prompts to avoid common mistakes.
-In particular, we emphasise direct and clear questions, with brief and specific answers to avoid speculation, repetition and ambiguity.
-Our example prompts are available in Appendix I.
-This stage results in 7,104 question-answer (QA) pairs that we use for the open-ended question-answering task.
-Automatic Distractor Generation.
-To construct the MCQ task, we use GPT-4 [42] to complement each question with 4 distractors.
-The distractors are crafted to provide plausible yet incorrect answers that fit within the movie’s context.
-To this end, our prompts are designed with specific criteria: distractors should be syntactically similar to the correct answer while being semantically different.
-Our distractors also incorporate diverse misdirections relative to movies (character confusion, plot adjustments), consider the movie genre and plausible scenarios, and include accurate but irrelevant information from other parts of the synopsis.
-This stage results in 7,104 Multiple-Choice Questions (MCQs), with 5 options per question with one correct and four incorrect answers.
-Manual QA Curation.
-To ensure the quality of questions and answers, we perform a thorough manual curation based on the following criteria.
-We remove questions that cannot be answered based on the synopsis, including factually incorrect, ambiguous, or subjective ones.
-Additionally, questions considered overly simplistic or answerable through external knowledge are excluded.
-This stage results in 4,885 MCQs from 1,078 movies.
-2.3 Dataset Analysis and Statistics
-Our SFD dataset contains a total of 1,078 unique films ranging in duration from 5 to 37 minutes and having an average length of 13 minutes (see Figure 6(a)).
-Overall, these movies amount to 243 hours and cover distinct genres, as illustrated in Figure 6(d).
-Each movie in the dataset is accompanied by a title, a logline which averages around 15 words, and a synopsis of about 97 words, equivalent to approximately 2 sentences, as shown in Figure 6(b).
-We use this metadata for question and answer generation as described in Section 2.2.
-The majority of films in SFD are in English and mainly originate from North America and Europe (see Figures 6(f) and 6(e)).
+Source and filtering.
+To build SF20K, we capitalize on the abundance of short films on video-sharing platforms.
+We target YouTube and Vimeo channels hosting high-quality and award-winning short films, some of which have been recognized by Oscars and BAFTA awards 111Among others, we use the YouTube channels Omeleto https://www.youtube.com/@Omeleto, Film Shortage https://www.youtube.com/@FilmShortage, and the Shortverse library https://www.shortverse.com.
+Videos, subtitles, and metadata are downloaded using yt-dlp 222https://github.com/yt-dlp/.
+We filter out videos less than three months old (before July 2024), and rely on platform safety filters to exclude sensitive, age-restricted, and private content.
+Annotation extraction.
+Since speech transcripts are available for only a portion of the dataset, we use Whisper Large-v3-Turbo [47] to generate missing subtitles.
+The metadata accompanying each short film includes the movie title, logline (a concise one-sentence plot summary), genre, release year, region of origin, and language.
+Additionally, we extract shot boundaries using PySceneDetect333https://github.com/Breakthrough/PySceneDetect, face tracks using DeepFace444https://github.com/serengil/deepface, and frame captions using Llama-3.2-Vision-11B [43].
+More details can be found in Appendix B.
+3.2 Scaling Up Instruction Tuning
+SF20K includes two tasks: Multiple-Choice Question Answering (MCQA) and Open-Ended Question Answering (OEQA).
+For both tasks, we use movie titles, loglines, and synopses — manually curated for the test set and automatically extracted for the training set (Section 3.2.1).
+Given this data, we automatically generate question-answer pairs by prompting GPT-4 [45] (Section 3.2.2).
+For SF20K-Test, we further refine the QA pairs through manual curation to ensure their relevance and correctness, and we augment this test set with additional manually created QA pairs (Section 3.2.3).
+We describe the generation process in detail below.
+The prompts used for question generation are available in Appendix H.
+3.2.1 Synopsis Preparation
+Human-written synopses for SF20K-Test.
+Among the 20K short films in SF20K, 1,072 come with detailed, human-written synopses.
+We use these high-quality synopses in the QA generation pipeline for the test set.
+Synthetic synopses for SF20K-Train.
+For the training set, we generate synthetic synopses by prompting GPT-4 [45] using subtitles.
+This process captures key story-level details, producing one synopsis per movie.
+3.2.2 Automatic Question Generation
+The automatic QA generation process is the same for both the train and test sets. See Appendix H for the exact prompts.
+QA pair generation.
+GPT-4 [45] is used to generate question-answer pairs based on movie titles, loglines, and synopses.
+Prompts are carefully designed to obtain concise, specific, and unambiguous questions, with brief answers.
+Distractor generation.
+To construct the MCQA task, we use GPT-4 [45] to complement each question-answer pair with four distractors.
+Distractors are crafted to be plausible yet incorrect, using syntactic similarity to the correct answer while maintaining semantic distinctions.
+Diverse misdirections, such as character confusion and irrelevant but accurate details, are included.
+3.2.3 Manual Curation
+SF20K-Train. The training set relies entirely on automatically generated QA pairs without manual curation.
+SF20K-Test. The test set undergoes manual curation to ensure high quality. Ambiguous, subjective, or overly simplistic questions are removed.
+| Dataset | Venue | Annotation | Average | Total | #QA Pairs | Multimodal | Long Term | Accessible | Unknown | Full | |
+| Duration (s) | Hours | Train | Test | To LLMs | Movies | ||||||
+| VideoQA datasets | |||||||||||
+| MSRVTT-QA [74] | ACM 2017 | Auto | 15 | 41 | 170,860 | 72,820 | ✗ | ✗ | ✓ | ✓ | - |
+| MSVD-QA [9] | ACM 2017 | Auto | 10 | 5 | 37,344 | 13,156 | ✗ | ✗ | ✓ | ✓ | - |
+| TGIF-QA [26] | CVPR 2017 | Auto | 3 | 47 | 139,414 | 25,751 | ✗ | ✗ | ✓ | ✓ | - |
+| ActivityNet-QA [81] | AAAI 2019 | Manual | 180 | 290 | 32,000 | 8,000 | ✗ | ✗ | ✓ | ✓ | - |
+| How2QA [55] | EMNLP 2020 | Manual | 60 | 366 | 35,205 | 4,400 | ✗ | ✗ | ✓ | ✓ | - |
+| NeXT-QA [71] | CVPR 2021 | Manual | 44 | 66 | 71,655 | 17,742 | ✗ | ✗ | ✓ | ✓ | - |
+| iVQA [38] | ICCV 2021 | Manual | 18 | 50 | - | 10,000 | ✗ | ✗ | ✓ | ✓ | - |
+| EgoSchema [41] | NeurIPS 2023 | Auto + Manual | 180 | 250 | - | 5,000 | ✗ | ✓ | ✓ | ✓ | - |
+| MovieQA datasets | |||||||||||
+| MovieQA [61] | CVPR 2016 | Manual | 203 | 380 | 4,318 | 1,258 | ✓ | ✓ | ✗ | ✗ | ✗ |
+| TVQA [31] | EMNLP 2018 | Manual | 76 | 460 | 122,036 | 15,253 | ✓ | ✗ | ✗ | ✓ | ✗ |
+| MovieChat [59] | CVPR 2024 | Manual | 459 | 180 | - | 14,000 | ✓ | ✓ | ✗ | ✓ | ✗ |
+| CinePile [49] | arXiv 2024 | Auto + Manual | 160 | 417 | 298,887 | 4,941 | ✓ | ✓ | ✓ | ✓ | ✗ |
+| SF20K (Ours) | - | Auto + Manual | 640 | 3,582 | 191,007 | 4,885 | ✓ | ✓ | ✓ | ✓ | ✓ |
+3.3 Dataset Statistics
+SF20K contains 20,143 short films ranging from 5 to 37 minutes, with an average length of 11 minutes per video (see Figure 5(a)), and totaling 3,582 hours.
+The dataset covers multiple genres, as illustrated in Figure 5.
 Examples of the dataset’s diversity can be seen in Figure 4.
-Additional data preprocessing and statistics are available in Appendix C, including clips, captions, location tags, and face tracks.
-In terms of video understanding tasks, SFD comprises 4,885 MCQs and OEQs, averaging 4.53 questions per film.
-Table 1 shows the comparison to both general video understanding (videoQA) and movie-oriented (movieQA) question-answering datasets.
-Notably, SFD exhibits the longest average video duration (821 seconds) among all datasets.
-Compared to videoQA datasets, SFD holds the properties of multimodality and long-term nature, similar to most movieQA datasets.
-Compared to other movieQA ones, SFD offers advantages in public availability of videos, data leakage (i.e., not exposed to LLMs), and story completeness (using full movies instead of trimmed clips, potentially impacting narrative coherence).
-Dataset | Venue | Annotation | Avg. | #QA Pairs | Multimodal | Long-Term | Accessible | Unknown | Full |
-Length (s) | To LLMs | Movies | |||||||
-General VideoQA datasets | |||||||||
-MSRVTT-QA [67] (test) | ACM 2017 | Auto | 15 | 72,820 | ✗ | ✗ | ✓ | ✓ | - |
-MSVD-QA [8] (test) | ACM 2017 | Auto | 10 | 13,156 | ✗ | ✗ | ✓ | ✓ | - |
-TGIF-QA [23] (test) | CVPR 2017 | Auto | 3 | 25,751 | ✗ | ✗ | ✓ | ✓ | - |
-ActivityNet-QA [75] (test) | AAAI 2019 | Manual | 180 | 8,000 | ✗ | ✗ | ✓ | ✓ | - |
-How2QA [50] (test) | EMNLP 2020 | Manual | 60 | 4,400 | ✗ | ✗ | ✓ | ✓ | - |
-NeXT-QA [65] (test) | CVPR 2021 | Manual | 44 | 9,178 | ✗ | ✗ | ✓ | ✓ | - |
-iVQA [35] | ICCV 2021 | Manual | 18 | 10,000 | ✗ | ✗ | ✓ | ✓ | - |
-EgoSchema [39] | NeurIPS 2023 | Manual + Auto | 180 | 5,000 | ✗ | ✓ | ✓ | ✓ | - |
-MovieQA datasets | |||||||||
-MovieQA [56] (test) | CVPR 2016 | Manual | 203 | 6,462 | ✓ | ✓ | ✗ | ✗ | ✗ |
-TVQA [29] (test) | EMNLP 2018 | Manual | 76 | 15,253 | ✓ | ✗ | ✗ | ✓ | ✗ |
-LVU [64] (test) | CVPR 2021 | Manual | 220 | 1,223 | ✓ | ✓ | ✓ | ✗ | ✗ |
-MovieChat [54] (test) | CVPR 2024 | Manual | 459 | 2,417 | ✓ | ✓ | ✗ | ✓ | ✗ |
-CinePile [45] (test) | arXiv 2024 | Manual + Auto | 160 | 4,940 | ✓ | ✓ | ✓ | ✓ | ✗ |
-SFD (Ours) | Manual + Auto | 821 | 4,885 | ✓ | ✓ | ✓ | ✓ | ✓ |
-Question Analysis.
-The generated question-answer pairs are long and diverse: the median answer length is 10 words and the vocabulary contains 8,928 different words.
-Furthermore, to ensure that the generated QAs cover various aspects of movies, we categorize them into four types (see Table 2):
+Most of the films are in English and originate mainly from North America and Europe (see Figure 5(d,e)).
+More samples can be seen in Appendix I.
+Annotations.
+SF20K includes 2.4M shots, with an average shot length of 5.52 seconds.
+Each shot has a 24-word caption describing settings and characters.
+Additional annotations include 1.1M subtitles and 1.08M face tracks.
+Tasks.
+SF20K-Train contains 191,007 questions, each suitable for both multiple-choice and open-ended formats, averaging 9.21 questions per movie.
+SF20K-Test includes 4,885 curated QA pairs, averaging 4.53 questions per movie.
+Comparison to other datasets.
+Table 1 compares SF20K to existing datasets.
+Notably, SF20K exhibits the longest average duration (640 seconds).
+Compared to video QA benchmarks, SF20K-Test distinguishes itself by its multimodal and long-term focus.
+Compared to movie QA benchmarks, SF20K-Test offers advantages in public availability of videos, limited data leakage (unlikely exposure to LLMs), and story completeness, as it uses full movies rather than trimmed clips, potentially impacting narrative coherence.
+3.4 Test Set Analysis
+This section analyzes SF20K-Test, focusing on question types and subset descriptions.
+Question types. SF20K-Test contains four question types:
+Setting-related questions address location and time of the movie, including geographical places, types of locations (e.g. home, office), historical periods, and times of day.
+Character-related questions focus on personal details (age, profession), traits, and relationships of the main characters.
+Story-related questions cover key events driving the plot.
+Theme-related questions explore the movie’s central idea.
+Test subsets. SF20K-Test comes with two specific subsets:
 1.
-Setting-related questions focus on the location and time of the movie.
-They may pertain to geographical locations (e.g., a specific city or country), a type of place (e.g., home, office), a historical period (e.g., a specific era), or a time of day (e.g., morning, night).
+SF20K-Test-Silent.
+This subset contains silent films with minimal or no dialogue, which accounts for around 10% of SF20K-Test.
+It consists of 90 movies (20.5 hours) and 419 questions.
+This subset is intended to challenge vision-language models that heavily rely on language information.
 2.
-Character-related questions focus on the individuals portrayed in the movie.
-They may concern personal information (age, gender, profession), traits, motivations, and relationships (siblings, friends).
-Note that character development, as part of the narrative, is excluded from this category.
-3. Story-related questions focus on the key sequence of events and interactions that compose the plot and narrative arc of the movie.
-4. Theme-related questions focus on the underlying message or central idea that the movie explores and conveys to its audience.
-Question Type | Questions | Answers |
-Setting | Where do John and Emma have their conversation? | On the bus. |
-Which city is the backdrop? | Liverpool. | |
-What decade does the movie take place in? | 1990s. | |
-Character | Who is the main character? | Josh. |
-What is the nationality of the soldier who becomes a prisoner? | German. | |
-How is Mr. Jones described? | As stern and harsh. | |
-Story | Can you name a specific tactic Tommy uses against Tiny? | He steals Tiny’s food at dinner and distracts her at a school-wide relay race. |
-What is the imminent threat to Greg’s residence? | The house is in foreclosure and the bank is about to repossess it. | |
-What triggers Carla to believe her husband Frank will surprise her for their anniversary? | A bouquet left on her doorstep. | |
-Theme | What is a major theme? | The theme revolves around reconnecting and the complexities of honesty between exes. |
-What main theme is explored through Sarah’s experience? | The theme of romantic expectations versus present realities. | |
-What is the main theme? | The processing of grief and reconciliation between siblings. |
+SF20K-Test-Expert.
+We ask human annotators to watch movies and craft questions, capturing complex elements like implicit information, key visual details essential to the story, and plot twists.
+This subset contains 50 movies (11.4 hours) and 570 questions, averaging 11.4 questions per movie (compared to 4.55 in SF20K-Test overall) and is designed exclusively for open-ended question answering.
 4 Experiments
+In this section, we introduce several experiments performed on the SF20K dataset.
+Section 4.1 examines data leakage in movie datasets.
+Section 4.2 assesses human performance on SF20K-Test.
+Section 4.3 benchmarks recent models across modalities.
+Section 4.4 reports performance on SF20K subsets.
+Section 4.5 evaluates temporal window effects on model accuracy.
+Finally, Section 4.6, analyzes the impact of instruction tuning on VLM performance.
 4.1 Data Leakage
-Modern LLMs, pre-trained on vast internet data, risk being exposed during training to information about commercial films, such as synopses, reviews, scripts, and transcripts.
+Modern LLMs, pre-trained on vast internet data, risk being exposed during pretraining to information about movies, such as synopses, reviews, scripts, and subtitles.
 This issue may result in models recalling answers directly without analyzing the video content, leading to biased benchmarks.
-In this section, we quantitatively assess the extent of data leakage on datasets.
-Specifically, we prompt LLMs to answer open-ended questions using only the movie title and compute the accuracy of responses, following [37].
-We experiment with 3 movie datasets: MovieQA [56], LVU [64] and our SFD, 5 open-sourced models: Gemma 2B [57], Mistral 7B/8x7B [25], LLaMA-3 8B/70B [58] and 4 commercial models Claude 3 Haiku/Sonnet, and GPT-3.5/4 [42].
-Figure 3 reports the results of the data leakage experiment and further details can be found in Appendix G.
-We observe that both MovieQA and LVU suffer from data leakage, reaching up to 71.3% and 76.0% accuracies.
-In contrast, thanks to the low presence of amateur film on the internet, SFD exhibits a maximum accuracy of 36.0%, indicating a low leakage issue.
-Furthermore, as expected, the extent of data leakage correlates with the knowledge level reflected by MMLU score [14] (see Appendix section G for more experiments), indicating that LLMs with more knowledge exhibit higher levels of memorization, leading to worse leakage issues.
-For instance, for both MovieQA and LVU, the zero-shot accuracy increases from approximately random level (20%) to more than 70% as the model size rises.
-Meanwhile, SFD maintains stable and low accuracies ranging between 19.7% and 36.0%, regardless of LLM knowledge variation, further indicating its low data leakage.
-This experiment reveals that relying solely on existing datasets to evaluate new methods is insufficient.
-Instead, our SFD offers a more objective and reliable test bed for long-term video understanding.
+We call this phenomenon data leakage.
+In this section, we quantitatively assess the extent of data leakage in movie datasets.
+For this purpose, we prompt LLMs to answer open-ended questions using only the movie title and compute the response’s correctness following [39].
+We experiment with three movie datasets: MovieQA [61], LVU [70], and SF20K-Test; five open-source models: Gemma 2B [62], Mistral 7B/8x7B [28], LLaMA-3 8B/70B [63]; and four commercial models: Claude 3 Haiku/Sonnet [4], GPT-3.5/4 [45].
+See Appendix C for more details.
+In Figure 3, we observe that both MovieQA and LVU suffer from data leakage, reaching up to 71.3% and 76.0% accuracies.
+In contrast, thanks to the low presence of amateur films on the internet, SF20K-Test exhibits a maximum accuracy of 36.0%, indicating a low leakage issue.
+More importantly, the extent of data leakage correlates with the knowledge level reflected by MMLU score [16], indicating that LLMs with more knowledge exhibit higher levels of memorization, leading to worse leakage issues.
+For instance, for both MovieQA and LVU, the zero-shot accuracy increases from approximately 20% to more than 70% as the LLM size rises.
+In contrast, SF20K-Test maintains stable and low accuracies ranging between 19.7% and 36.0%, regardless of LLM knowledge variations, further indicating limited knowledge of amateur movies.
+| SF20K-Test | MovieQA | |||
+| No Anonymization | 31.5 | 71.3 | ||
+| Movie Title Removed | 32.8 | +1.3 | 53.9 | -17.4 |
+| Full Anonymization | 28.1 | -3.4 | 31.4 | -39.9 |
+Anonymization and shuffling. To better understand the impact of movie titles on LLM performance, we perform two additional experiments.
+(i) Anonymization.
+We examine the zero-shot performance of GPT-4 in three settings: (i) providing movie titles, (ii) removing movie titles, and (iii) removing movie titles with additional question anonymization.
+Question anonymization means replacing character names in the questions with generic placeholders, as this may help LLMs to recall specific movies (e.g., Harry Potter, James Bond).
+It enables us to measure relative accuracies, i.e.
+the performance improvement of an LLM if it can guess the movie.
+For MovieQA, removing movie titles causes a significant drop in performance (-17.4%), with further reduction after full anonymization (-39.9%), as shown in Table 2.
+In sharp contrast, the performance on the SF20K-Test remains overall stable.
+This indicates that the ability to identify the movie is the main reason why LLMs perform so well in commercial movies in our data leakage experiment, while this does not apply to amateur films.
+(ii) Shuffling.
+We examine the data leakage performance of Llama-3-70B in a scenario where movie titles are replaced with random titles from the same dataset instead of providing the correct ones.
+As shown in Table 3, this substitution causes a substantial drop in accuracy for MovieQA, even falling below the fully anonymized setting.
+This suggests that LLMs rely heavily on recognizing commercial movie titles for their predictions.
+On the contrary, for SF20K-Test, replacing correct titles with random ones does not notably affect performance, likely because the LLM does not recognize these amateur films in either case.
+These experiments show that (i) evaluating new methods on commercial movies can be misleading, and (ii) SF20K-Test provides a more reliable benchmark for long-term video understanding.
+| SF20K-Test | MovieQA | |||
+| Correct movie title | 24.1 | 51.9 | ||
+| Random movie title | 19.2 | -4.9 | 10.9 | -41.0 |
+4.2 User study
+To verify the answerability of our multiple-choice questions and assess the upper limit of SF20K-Test, we conduct three user studies according to the following modality definitions: (1) Vision-Language (VL), i.e.
+full video with audio and subtitles; (2) Vision-only (V), i.e.
+muted videos; (3) Language-only (L), i.e.
+plain text subtitles.
+For each question, all participants are asked to select the correct answer (see Appendix D for more details).
+Table 4 (last row) reports the results.
+We observe that when provided with the full multimodal information, participants answer questions with high accuracy (89.8%).
+As expected, removing modalities lowers accuracy.
+Specifically, when using only subtitles, the performance drops to 70.9%, whereas the vision-only performance decreases to 59.0%.
+4.3 Baselines
+In this section, we evaluate recent open-source and commercial vision-language models on SF20K-Test.
+In particular, we benchmark eleven open-source models: FrozenBiLM [6], mPLUG-Owl2 [80], TimeChat [52], Video-Llava [36], MovieChat [59], Llava-OneVision [32], Long-Llava [67], LongVA [84], Llava-Next-Video [85], Llama-3.2-Vision [43], and Pixtral [3]; and two commercial models: LloVi [83] (based on GPT-3.5 [45]), and GPT-4o-mini [46].
+See Appendix E for more details on prompts and metrics.
+The models are tested in a zero-shot video question-answering setting, adapted for long-form video understanding by incorporating subtitles and sampling the maximum number of frames possible to fit in a single A100 GPU.
+For multiple-choice questions, we compute the accuracy score by extracting the option letter choice from the model’s response.
+For open-ended questions, we rely on GPT-3.5 to compute the similarity between the predicted and correct answers, as detailed in [39].
+All methods are evaluated across three modalities to better assess the contribution of each: Vision-Only (V) with video frames, Language-Only (L) with subtitles, and Vision-Language (VL) combining both.
+The results are summarized in Table 4.
+| Method | Multiple-Choice QA | Open-Ended QA | ||||
+| V | L | VL | V | L | VL | |
+| Random | 20.0 | 20.0 | 20.0 | - | - | - |
+| Open-source models | ||||||
+| FrozenBiLM [79] | 23.4 | 38.2 | 38.6 | - | - | - |
+| mPLUG-Owl2 [80] | 38.3 | 20.7 | 21.3 | 22.1 | 1.8 | 1.6 |
+| TimeChat [52] | 25.5 | 6.4 | 31.8 | 26.4 | 9.4 | 5.9 |
+| Video-Llava-7B [35] | 32.4 | 21.3 | 45.7 | 8.4 | 10.6 | 8.2 |
+| MovieChat [59] | 8.4 | 6.4 | 8.0 | 14.0 | 15.7 | 11.8 |
+| Llava-OneVision-0.5B [32] | 29.2 | 35.4 | 37.8 | 14.3 | 20.2 | 22.6 |
+| Long-Llava-7B [67] | 52.1 | 70.4 | 72.6 | 17.4 | 36.7 | 33.5 |
+| LongVA-7B [84] | 58.5 | 73.8 | 70.0 | 26.3 | 51.7 | 36.6 |
+| Llava-OneVision-7B [32] | 63.2 | 74.7 | 78.2 | 25.1 | 35.3 | 37.7 |
+| Llava-Next-Video-7B [85] | 32.8 | 52.3 | 64.2 | 27.4 | 42.2 | 42.7 |
+| Llama-3.2-Vision-11B [43] | 57.9 | 72.3 | 75.0 | 24.4 | 45.2 | 47.8 |
+| Pixtral-12B [3] | 38.5 | - | 74.4 | 29.5 | 49.4 | 49.8 |
+| Commercial models | ||||||
+| LloVi [83] (GPT-3.5-based) | 30.8 | 64.2 | 55.6 | 16.2 | 40.3 | 24.7 |
+| GPT-4o-mini [46] | - | - | 79.4* | - | - | 60.0* |
+| Human | 59.0 | 70.9 | 89.8 | - | - | - |
+∗ indicates that inference is done on only 10% of the test set to limit costs.
+Multiple-Choice Question Answering (MCQA).
+Our results reveal that several models exhibit strong performance, with GPT-4o-mini achieving the highest accuracy of 79.4%.
+Notably, recent open-source models are closing the gap with commercial ones; for example, Llava-OneVision-7B lags behind by just 1.2 percentage points.
+However, there remains significant room for improvement, as human performance reaches 89.8%.
+Open-Ended Question Answering (OEQA).
+This task is more challenging, as evidenced by significantly lower accuracy across all models, ranging from 1.6% to 60.0%.
+GPT-4o-mini again leads in performance.
+On the other hand, open-source models face a larger performance gap.
+For instance, Pixtral-12B achieves a maximum accuracy of only 49.8%.
+Appendix F provides an analysis of failure cases (see Table 12).
+Modality Ablations.
+For both tasks, recent models demonstrate improved modality integration, as multimodal performance consistently surpasses unimodal performance.
+However, language remains the dominant modality, as evidenced by higher accuracy in language-only settings compared to vision-only settings, and a relatively small performance gap between language-only and multimodal settings (+19.9% and +0.4%, respectively, for Pixtral-12B on open-ended questions).
+Overall, our results indicate that (1) recent models perform reasonably well on multiple-choice questions, though there remains substantial room for improvement compared to human performance; (2) open-ended question answering poses a much greater challenge, with all models struggling, particularly open-source models who lag behind their commercial counterparts; and (3) performance relies heavily on subtitles, suggesting that better integration of the visual modality could improve accuracy.
+| Model | SF20K-Test | Silent | SF20K-Test-Expert | ||||
+| V | L | VL | V | V | L | VL | |
+| Open-source models | |||||||
+| Video-LLaVA-7B [35] | 8.4 | 10.6 | 8.2 | 7.2 | 9.0 | 5.0 | 3.5 |
+| Llava-Next-Video-7B [85] | 27.4 | 42.2 | 42.7 | - | 14.2 | 10.3 | - |
+| Llava-OneVision-0.5B [32] | 14.3 | 20.2 | 22.6 | 17.6 | 12.5 | 12.0 | 19.5 |
+| Llava-OneVision-7B [32] | 25.1 | 35.3 | 37.7 | 29.0 | 21.5 | 20.8 | 29.0 |
+| Long-Llava-7B [84] | 17.4 | 36.7 | 33.5 | 20.9 | 17.2 | 22.0 | 29.5 |
+| Llama-3.2-Vision-11B [43] | 24.4 | 45.2 | 47.8 | 29.0 | 17.2 | 29.8 | 31.0 |
+| Pixtral-12B [3] | 29.5 | 49.4 | 49.8 | 17.0 | 17.0 | 30.9 | 33.0 |
+| Commercial models | |||||||
+| GPT-4o-mini [2] | - | - | 60.0* | - | - | - | - |
+∗ indicates that inference is done on only 10% of the test set to limit costs.
+4.4 Performance on SF20K-Test subsets
+In this section, we evaluate recent VLMs on two subsets: SF20K-Test-Silent (silent movies, without dialogs) and SF20K-Test-Expert (manual questions created by expert annotators).
+Both subsets are evaluated in an OEQA setting.
+Table 5 reports the results.
+For silent movies, the modality ablation is not performed since there are no transcripts available.
+SF20K-Test-Silent.
+Interestingly, all models present a significant performance drop in the silent subset compared to the normal subset, with accuracy not exceeding 30%.
+Note that these questions are similar in type and difficulty to those in the full SF20K-Test, where models reach 49.8% accuracy.
+For example, Llama-3.2-Vision-11B performance declines by 18.8%, and other models show similar trends.
+Moreover, performance differences between models are reduced in this subset, with both Llava-OneVision-7B and Llama-3.2-Vision-11B now ranking as top-performing models.
+This consistent drop in accuracy suggests that models are not effectively leveraging visual information alone.
+SF20K-Test-Expert.
+On challenging questions, all models experience a performance drop, with both Pixtral-12B and Llama-3.2-Vision-11B showing a 16.8% decline.
+Moreover, we observe that the Language-only (L) and Vision-Language (VL) performances are similar, suggesting that subtitles (L) continue to be the primary contributor to model performance, which indicates the models’ limited reliance on visual information.
+| Fine-tuned | Multiple-Choice QA | Open-Ended QA | ||||||||||
+| V | L | VL | V | L | VL | |||||||
+| ✗ | 29.2 | 35.4 | 37.8 | 14.3 | 20.2 | 22.6 | ||||||
+| ✓ | 45.7 | +16.5 | 61.2 | +25.8 | 63.0 | +25.2 | 17.8 | +3.5 | 22.5 | +2.3 | 26.6 | +4.0 |
 4.5 Temporal Window Study
-To confirm that our tasks require long-form video understanding, we conduct a temporal window study with benchmarks at three levels: shot-, scene-, and movie-level.
-A shot, defined as a continuous video clip, typically amounts to around 151 per movie.
-At the Shot-Level, the model uses data from a single shot, including partial subtitles and visual data, with predictions aggregated by taking the maximum logits.
-Scene-Level inference aggregates data from approximately 10 shots in a similar manner.
-Movie-Level inference utilizes the entire film’s data, as detailed in Section 3.3.
-To further analyze the behaviors of each modality, we ablate at each temporal level three combinations of modalities: Vision, Language, and multimodal Vision-Language.
-This study is conducted with two top-performing methods from Section 3.3: FrozenBiLM and LLoVi.
-Further details on the experiment can be found in Appendix E.
+To confirm that our tasks require long-form video understanding, we conduct a temporal window study by performing inference at three levels: shot-, scene-, and movie-level.
+Shots, defined as continuous video clips, last 5.52 seconds on average.
+At the Shot-Level, the model uses data from one shot, including partial subtitles and visual data, with predictions aggregated by taking the maximum logits over all shots.
+Scene-Level inference aggregates data from approx.
+10 shots, in a similar manner.
+Movie-Level inference utilizes the entire film’s data.
+To further analyze the behavior of each modality, we ablate at each temporal level three combinations of modalities: Vision, Language, and Vision-Language.
+This study is conducted with two methods: FrozenBiLM and LLoVi, for the MCQA task.
+More details in Appendix G.
 Our results, shown in Figure 6, reveal the following trends: (1) Both Language-only and Vision-Language settings show substantial improvements with larger temporal windows.
@@
 The multimodal setting of LLoVi’s accuracy increases from 50.1% to 55.6% at the scene-level but shows only marginal gains at the movie-level.
-This plateau may be attributed to LLoVi’s naive fusion approach, which combines visual captions with subtitles, limiting further improvements; (2) The Vision-only modality exhibits low and stagnant performance across all levels suggesting that the current handling of visual data is inadequate for this task.
-Additional results for other methods are presented in Appendix E.
-In conclusion, broader temporal windows significantly enhance SFD task performance, especially in language-only settings, underscoring the importance of long-form understanding in our design.
-3.3 Baseline Comparison
-In this section, we benchmark recent videoQA methods on SFD and evaluate 7 open-source models in three categories: FrozenBiLM [5], mPLUG-Owl2 [74] and Video-LLaVA [33] are VLMs that project visual frames into embedding of an LLM with the help of adapters; MovieChat [54] and TimeChat [47] have more advanced and specific settings, the former introduces a memory mechanism to handle long sequences and the latter proposes a timestamp-aware frame encoder for time-sensitive tasks; LLoVi [79] and LangRepo [26] are text-based methods: they both rely on visual captions on top of an LLM.
-LangRepo proposes an extra summarization and chain-of-thoughts [62] reasoning prompting.
-These models are tested in a zero-shot video question-answering setting, adapted for long-form video understanding by incorporating subtitles and sampling the maximum number of frames possible.
-In the Multiple-Choice Question answering (MCQ) benchmark, we calculate the accuracy score based on the logits of the predicted options where applicable [5, 33, 74].
-For models that generate plain-text answers, we evaluate them using a method similar to the Open-Ended Question answering (OEQ) where we rely on GPT-3.5 to compute the similarity between the predicted and correct answers, as detailed in [37].
-All methods are evaluated across three modalities to better assess the contribution of each: Vision-Only (V) with video frames, Language-Only (L) with subtitles, and Vision-Language (VL) combining both.
-See Appendix D for more details.
-The results are summarized in Table 3.
-User Study.
-To verify the answerability of our multiple-choice questions and assess the upper limit of SFD, we conducted three user studies accordingly to the aforementioned modality definitions: (1) Vision-Language (VL),—full video with audio and subtitles; (2) Vision-only (V)-muted videos; (3) Language-only (L)-plain text subtitles.
-For each question, all participants were asked to select the correct answer.
-Table 3 (last row) reports the results.
-More details on the user study are available in Appendix F.
-We observe that when provided with the full multimodal information, participants answer questions with high accuracy (89.8%).
-As expected, removing modalities lowers accuracy.
-Specifically, when using only subtitles the performance is 70.9%, whereas the vision-only performance drops to 59.0%.
-Method | Venue | % Accuracy | |||||
-Multiple-Choice QA | Open-Ended QA | ||||||
-V | L | VL | V | L | VL | ||
-Random | 20.0 | 20.0 | 20.0 | - | - | - | |
-FrozenBiLM [5] | NeurIPS 2021 | 23.4 | 38.2 | 38.6 | - | - | - |
-mPLUG-Owl2 [74] | CVPR 2024 | 38.3 | 20.7 | 21.3 | 22.1 | 1.8 | 1.6 |
-Video-LLaVA [33] | arXiv 2023 | 34.2 | 21.3 | 24.7 | 19.2 | 6.4 | 8.0 |
-LLoVi [79] | arXiv 2023 | 30.8 | 64.2 | 55.6 | 16.2 | 40.3 | 24.7 |
-LangRepo [26] | arXiv 2024 | 29.0 | 32.1 | 31.0 | 3.5 | 10.4 | 9.5 |
-MovieChat [54] | CVPR 2024 | 8.4 | 6.4 | 8.0 | 14.0 | 15.7 | 11.8 |
-TimeChat [47] | CVPR 2024 | 25.5 | 6.4 | 31.8 | 26.4 | 9.4 | 5.9 |
-Human | 59.0 | 70.9 | 89.8 | - | - | - |
-Multiple-Choice Question answering (MCQ).
-These results highlight significant insights.
-Overall, model performance is generally poor, with LLoVi being a notable exception in the language-only setting (64.2%), driven largely by subtitles.
-Some models, such as FrozenBiLM, perform adequately in the vision-only setting (38.3%), but their performance drops when combining both modalities, indicating integration issues.
-Compared to human performance, there is still substantial room for improvement.
-Accuracy is on par in the language-only setting, with only a 6.7% difference.
-However, as the best models mainly rely on subtitles, there is a significant gap in the multi-modal setting, where human performance reaches 89.8%.
-This discrepancy can be explained by the vision-only setting, where the gap reaches 20.7%.
-These results highlight that while models perform reasonably well with textual data, they struggle to effectively integrate visual information, underscoring the need for advancements in multi-modal fusion techniques to bridge the gap between human and model performance.
-Open-Ended Question answering (OEQ).
-This task proves even more challenging, with lower accuracy across all models, ranging from 3.5% to 40.3%.
-Subtitles-only models, particularly LLoVi, lead in performance, highlighting the dominance of language processing in understanding and answering complex questions.
-These findings emphasize the difficulty of the task and the need for improved methods to handle and integrate multi-modal data effectively.
-Overall, our results suggest that (1) text (in the form of subtitles) is a stronger cue than visual information for movie question-answering; (2) modern multimodal methods (max accuracy at 60.0%, see Table 3) fall behind human evaluation; and more importantly (3) there is a large room for improvement in the visual aspect, where the gap in performance between modern methods (average accuracy 38.3%) and user study (59.0% accuracy) is still very high.
-4 Related Work
-VideoQA benchmarks.
-Questions in such benchmarks can be framed to assess the reasoning, memory, and comprehension skills of both humans and algorithms.
-Hence, many datasets have been made available through the years, covering visual descriptions [69, 8], temporal action reasoning [65], compositional reasoning [19], social intelligence [76], instructional videos [71], egocentric videos [39, 24, 34], movies [56, 29, 38], among others [23, 67, 23, 32, 75, 63].
-However, these datasets feature short videos and typically require reasoning on only a few frames to solve the task at hand, as shown by temporal certificates in [39].
-VideoQA methods.
-To address video question answering, most modern approaches are based on large-scale pre-trained models, by deploying either multimodal contrastive learning [15, 31, 61, 78, 77, 10, 9] or visually-conditioned large language models [30, 11, 68, 74, 2, 70, 73].
-For instance, mPLUG-Owl2 [68] enhances LLMs with visual capabilities, while FrozenBiLM [73], a frozen bidirectional language model, uses lightweight adapters and masked language modelling to excel in zero-shot tasks.
-FrozenBiLM excels in zero-shot video question-answering without using explicit supervision.
-More recently, Video-LLaVA [33] used a simple projection layer to connect a visual encoder and a large language model, fine-tuned with visual instruction tuning data.
-Long-form video understanding.
-As new long-context models emerge [46], we need more complex benchmarks.
-Long-form video understanding evaluates the long-term reasoning capabilities of models.
-Some datasets have been introduced in recent years, mostly for egocentric videos [39] and movies [64, 53, 21, 56].
-Instead, LVU [64] involves long movie clips but the subsequent tasks are relatively limited, focusing on irrelevant aspects like ’like’ ratio and view count prediction.
-[39] proposes multiple-choice questions on 3-minute videos that require on average 100 seconds of video watching to be answered correctly.
-[53, 21] introduce an audio description task, which is currently tackled at the clip level but could be extended over the whole movie.
-Finally, closer to our work, [56] and [54] propose a movie question-answering datasets focused on movie plots, settings and characters.
-However, they do not release any video data because of copyright issues.
+This plateau may be attributed to LLoVi’s naive fusion approach, which combines visual captions with subtitles, limiting further improvements; (2) The Vision-only modality exhibits low and stagnant performance across all levels, suggesting that the current handling of visual data is inadequate for this task.
+In conclusion, larger temporal windows substantially enhance the performance on SF20K-Test, especially in language-only setting.
+This underscores the long-term nature of our proposed story-level video tasks.
+4.6 Instruction tuning
+In this experiment, we fine-tune Llava-OneVision-0.5B [32] on SF20K-Train (Section 3.2), using LoRA for efficiency.
+The input includes video frames, subtitles, and the question.
+We use next-token prediction to generate responses.
+In the multiple-choice case, we add multiple-choice options to the query and predict the correct option index in addition to the answer.
+Fine-tuning is performed over 191,007 samples from 19,676 movies for 5 epochs, with a learning rate of 1.0e-4, 50 warmup steps, batch size 8, gradient norm 1.0, and LoRA parameters (r=8, alpha=8, dropout=0.1).
+Samples are randomly presented as multiple-choice or open-ended questions.
+Instruction tuning performance.
+Table 6 reports the results, showing that instruction tuning consistently enhances performance across all modalities and tasks compared to the pre-trained model.
+For multiple-choice QA, performance improves significantly by 25.2%, reducing the gap with its 7B counterparts.
+In the open-ended QA setting, while the improvement is more modest (+4.0%), it remains meaningful.
+Notably, the vision-only modality also benefits from instruction tuning, despite the instruction data being generated from subtitles.
+Effect of the training size.
+Figure 9 shows the performance of Llava-OneVision-0.5B [32] when instruction-tuned for multiple-choice question answering on SF20K-Train, with varying numbers of training videos.
+We subsample the full dataset at incremental values.
+Results indicate a logarithmic improvement in accuracy as the training set size increases, rising from 46.5% to 63.0%.
+Results are displayed in log scale.
+These findings highlight the scalability benefits of instruction tuning on SF20K-Train.
 5 Discussion
-We introduced the Short Film Dataset (SFD), a long-form video understanding benchmark featuring complete and narrative-driven amateur short movies.
-It includes two types of QA tasks, multiple-choice and open-ended, that are automatically generated by LLMs and manually curated by human annotators.
-Compared to other long-form video datasets, SFD stands out by offering richer, story-oriented content with tasks specifically tailored for long-term reasoning, as validated by our experiments.
-Unlike most movie datasets that use copyrighted commercial films prone to data leakage with LLMs, SFD relies on amateur films, which are publicly accessible and have a limited online presence.
-Furthermore, our analysis indicates that while language-based methods achieve performance levels comparable to humans, state-of-the-art vision-based and multimodal methods fall behind human evaluation and they still have considerable room for improvement, highlighting the need for further advancements.
-In conclusion, we believe that our SFD paves the way for accessible, and comprehensive long-term movie understanding that is not known a priori by LLMs, thus helping the community develop robust methods.
-Our dataset and code will be made publicly available.
-6 Acknowledgement
+We introduced Short-Films 20K (SF20K), a long-form video understanding dataset featuring 20,143 amateur movies.
+It includes two QA tasks, multiple-choice and open-ended, automatically generated by LLMs and, in the test set, manually curated by human annotators.
+Compared to other video datasets, SF20K stands out by offering richer, story-oriented content with tasks specifically tailored for long-term reasoning, as validated by our experiments.
+Unlike other movie datasets that use copyrighted commercial movies prone to data leakage with LLMs, SF20K relies on amateur films, which are publicly accessible and have limited online presence.
+Our analysis indicates that state-of-the-art multimodal methods fall behind human accuracy, with performances being mostly driven by the language modality (i.e.
+subtitles), implying rooms for improvement and the need for further advancement of vision-language models.
+Finally, we show that instruction-tuning VLMs on SF20K-Train improves performance across modalities and tasks.
+In conclusion, we believe that SF20K paves the way for accessible and comprehensive movie understanding that is not known a priori by LLMs, thus helping the community to develop robust methods.
+6 Acknowledgments
 This work was supported by ANR-22-CE23-0007, Hi!Paris grant and fellowship, and was granted access to the High-Performance Computing (HPC) resources of IDRIS under the allocations 2023-AD011014489 made by GENCI.
-We would like to thank Nicolas Dufour, Robin Courant, Pierre Vassal, and the anonymous reviewers for their insightful comments and suggestions.
+We would like to thank Nicolas Dufour, Robin Courant, and Pierre Vassal for their insightful comments and suggestions.
 We also express our gratitude to the annotators of the human study for their help.
 Appendix
-In this appendix, we show additional materials of the paper Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding. We present the following items:
-Appendix A Checklist
-1.
-For all authors…
-Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
-Did you describe the limitations of your work? [Yes]
-Did you discuss any potential negative societal impacts of your work? [Yes] See Supplementary Materials Section B
-Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Supplementary Materials Section B
-2.
-If you are including theoretical results…
-Did you state the full set of assumptions of all theoretical results? [N/A]
-Did you include complete proofs of all theoretical results? [N/A]
-3.
-If you ran experiments (e.g. for benchmarks)…
-Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://huggingface.co/datasets/rghermi/sfd
-- (b)
-Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
-[No] Our experiments consist of benchmarking various aspects of our dataset.
-For a given compute, instead of performing one experiment several times (with different random seeds), we chose to experiment with several different multimodal methods and multiple LLMs.
-We observed that they all report similar performances and overall follow the same trend; hence, we argue that our findings are generalizable.
-Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 3.3
-4.
-If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
-If your work uses existing assets, did you cite the creators? [Yes] See Section B.
-Did you mention the license of the assets? [Yes] See Section B.
-Did you include any new assets either in the supplemental material or as a URL? [N/A]
-Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Section B.
-Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
-5.
-If you used crowdsourcing or conducted research with human subjects…
-Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section F.
-Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
-Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
-Appendix B Ethical considerations
+In the following, we present the following items:
+Ethical considerations
+Data pre-processing and statistics
+Technical details on the data leakage experiment
+Human study
+Technical details on baselines
+Failure cases analysis
+Technical details on the temporal window study
+Prompts
+Samples
+Datasheet for datasets
+Appendix A Ethical considerations
 1.
 Copyright and License:
-We do not distribute raw video content; instead, we provide only URLs redirecting to YouTube, where the full copyright and licensing rights of creators are acknowledged.
-Additionally, by sharing the URLs, YouTube’s Terms of Service regarding exclusive rights are respected.333https://www.youtube.com/static?template=terms (Section: License to Other Users).
-The dataset will be used solely for academic research purposes.
-To further protect the metadata and other information, we apply the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license to our dataset.
+We do not distribute raw video content; instead, we provide only URLs redirecting to YouTube and Vimeo, where the full copyright and licensing rights of creators are acknowledged.
+Additionally, by sharing the URLs, YouTube’s and Vimeo’s Terms of Service regarding exclusive rights are respected.222https://www.youtube.com/static?template=terms (Section: License to Other Users), https://vimeo.com/terms (Section: Licenses Granted by You).
+The dataset is intended exclusively for academic research purposes.
+To ensure the protection of the dataset and all its associated elements, we use the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International (CC BY-NC-SA 4.0) license.
 2.
@@
 Bias: Collected movies are primarily from North America and Europe, mainly English-speaking, the under-representation of the content towards non-English-speaking population and culture, potentially leading to unbalanced, biased video understanding model training and testing.
-Impact on creative process: Progress on movie understanding may encourage filmmakers to tailor their work to optimize for what is favoured by algorithms, potentially stifling creative freedom and depriving innovation.
-Appendix C Data Preprocessing and Statistics
-SFD | |
-#Movies | 1,078 |
-#Clips | 161,321 |
-#Clips / Movie | 151 |
-Avg. Clip Len. (s) | 5.43 |
-#Captions | 161,321 |
-#Location Tags | 161,321 |
-SFD | |
-#Face Instances | 56,265 |
-#Unique Faces | 2,598 |
-#Unique Faces / Movie | 2.41 |
-#Subtitles | 23,820 |
-SFD, alongside the videos and subtitles (see Section 2), also offers a comprehensive set of annotations: clips (shots), shot captions and scales, location tags, and face tracks.
-These resources, together with tools for downloading, preprocessing, and annotation, will be made available as part of the dataset release.
-Below, we detail their creation processes.
-Clips (Shots).
-In the context of movies, a clip corresponds to a shot, characterized by its composition, length, and camera movement.
-We use PySceneDetect444https://www.scenedetect.com/ to segment movies in SFD into 161,321 shots, with each movie containing an average of 151 shots, as shown in Table 5(a).
-The number of shots varies by genre; for example, comedies average 163 shots, while dramas have about 146.
-Clip information.
-For each shot, we extract detailed shot-level information.
-We generate shot captions using image and video captioning tools (LLaVA [36], BLIP-2 [30], Video-LLaVA [33]), and we identify locations using CLIP features [43] based on a curated list of locations from MovieCLIP [6].
+Impact on creative process: Progress on movie understanding may encourage filmmakers to tailor their work to optimize for what is favoured by algorithms, potentially limiting creative freedom and depriving innovation.
+Appendix B Data pre-processing and statistics
+SF20K provides not only videos but also a comprehensive set of annotations to facilitate deeper analysis and understanding.
+These annotations include subtitles, shot boundaries, keyframe captions, and face tracks.
+Together with tools for downloading, pre-processing, and annotating the data, these resources are released as part of the dataset.
+Below, we detail the processes involved in creating these annotations.
+| #Movies | #Hours | #QA Pairs | Source | Annotation | |
+| SF20K | 20,143 | 3,584 | 191,007 | Subtitles | Auto |
+| SF20K-Test | 1,072 | 244 | 4,885 | Synopsis | Auto + Manual |
+| SF20K-Test-Silent | 90 | 20 | 419 | Synopsis | Auto + Manual |
+| SF20K-Test-Expert | 50 | 11 | 570 | Movie | Manual |
+| SF20K | |
+| #Movies | 20,143 |
+| #Shots | 2,396,583 |
+| #Shots / Movie | 119 |
+| Avg. Shot Len. (s) | 5.52 |
+| #Shot Captions | 2,396,583 |
+| #Subtitles | 1,088,140 |
+| #Face Tracks | 1,083,850 |
+Shot boundaries.
+In the context of movies, a shot is a continuous sequence of frames captured without interruption by a single camera, often representing a single perspective or action within the scene.
+This corresponds to a clip in other video datasets.
+We use PySceneDetect333https://www.scenedetect.com/ with a threshold equal to 3.0 to segment the 20,143 movies in SF20K into 2,396,583 shots, with each movie containing an average of 119 shots.
+Keyframe captions.
+For each shot, a detailed caption is generated by extracting the middle frame of the shot and processing it with Llama-3.2-Vision [43].
+The model is specifically prompted to identify and describe relevant details, including the time, location, character appearances, mood, and overall atmosphere.
+We end up with one caption per shot, i.e.
+2,396,583 captions, for a total of 59,760,112 words.
+These textual annotations provide a good basis for further work on video generation models.
+Subtitles. We extract missing subtitles from the audio with Whisper Large-v3 Turbo [47], using the translation mode. We end up with 1,088,140 dialog sentences, for a total of 12,038,891 words.
 Face tracks.
-We use an advanced facial recognition pipeline 555https://github.com/serengil/deepface to identify over 56k+ face instances in the dataset, corresponding to 2.6k+ unique characters, with an average of 2.41 characters per movie, as shown in Table 5(b).
-These annotations enrich the dataset, providing deeper insights for tasks related to character analysis and interaction within the films.
+We use a facial recognition pipeline 444https://github.com/serengil/deepface to identify over 1,083,850 face tracks.
+We first detect faces with RetinaFace [14], and then compute face tracks by matching face detections among consecutive frames of the same shot with IoU matching.
+These annotations enrich the dataset, providing deeper insights for tasks related to character analysis and interaction within films.
+Word cloud analysis.
 In Figure 8, we perform a word cloud analysis for three video domains commonly used in the literature: instructional videos [77], egocentric videos [41], and movies (ours).
@@
 home).
-These clearly show the story component of SFD.
-Appendix D Technical details on baselines
-In this section, we present additional details and results for the baselines discussed in Section 3.3 and Table 3 main paper, where we report the benchmark performance of the state-of-the-art video understanding methods on our SFD.
-FrozenBiLM [5].
-We use FrozenBiLM pre-trained on WebVid10M[5].
-Following[39], we reformulate the question-answering task for this model with the following prompt: ‘Question: {question} Is it {option_i}?’.
-We get logits for the yes/no tokens for each option and take the maximum as the final prediction.
-We report results with and without subtitles.
-We compute the accuracy score, by comparing the predicted option index and the correct option index.
-As it is a masked language model, FrozenBiLM is not suited to open-ended question answering.
-mPLUG-Owl2 [74].
-We use the instruction-tuned version (‘MAGAer13/mplug-owl2-llama2-7b’), based on ViT-L and LLaMa-7b.
-Inference is the same as FrozenBiLM.
-In the open-ended question-answering setting, we only provide the question and record a free-form text output: ‘Question: {question} Please answer with a short sentence.’.
-Following [37], we use GPT-3.5 (‘gpt-3.5-turbo’) to compute the similarity between the predicted and the correct answer.
-We report this LLM-based accuracy.
-Video-LLaVA [33].
-We use the fine-tuned version (‘LanguageBind/Video-LLaVA-7B-hf’), based on LanguageBind and Vicuna-13B v1.5.
-Inference is the same as mPLUG-Owl2, in both multiple-choice and open-ended question-answering settings.
-LLoVi [79].
-Following the official implementation, we use LLaVA (‘llava-hf/llava-1.5-13b-hf’) to extract captions at 0.5 FPS and GPT-3.5 (‘gpt-3.5-turbo’) as a backbone LLM.
-We experiment with the prompting strategy proposed in the original paper [79].
-The prompt for multiple-choice question answering is as follows:
-‘Please provide a single-letter answer (A, B, C, D, E) to the following multiple-choice question, and your answer must be one of the letters (A, B, C, D, or E).
-You must not provide any other response or explanation.
-You are given some language descriptions of a video.
-The video is {duration} seconds long.
-Each sentence describes a {clip_length}s clip.
-The descriptions are sequential and non-overlapping which cover the whole video exactly.
-Here are the descriptions: {captions}.
-You are going to answer a multiple-choice question based on the descriptions, and your answer should be a single letter chosen from the choices.
-Here is the question: {question}.
-Here are the choices.
-A: {option_0} B: {option_1} C: {option_2} D: {option_3} E: {option_4}")’
-LangRepo [26].
-Similar to LLoVi, LangRepo is another LLM-backed video understanding method.
-We use the open-sourced official implementation, Mixtral 87B (‘mistralai/Mixtral-8x7B-Instruct-v0.1’).
-Unlike LLoVi, LangRepo operates in two stages: The first is summarization, where LangRepo consolidates similar information from descriptions, captions, subtitles, or both, using specific prompts following the official implementation:
-‘You are a helpful expert in video analysis.
-You are given a list of {num_to_rephrase} language descriptions for a video.
-Each sentence describes a {clip_length}s clip.
-Here are the descriptions as a list:{memory}.
-Please summarize and rephrase each item in the list as a single sentence of {num_words_in_rephrase} words.
-Keep the same original subject.
-Keep all information intact without leaving anything out.
-Return only the rephrased list of {num_to_rephrase} descriptions in the same order, without additional details.’
-In the second stage, the question-answering is based on the rephrased summarization with prompts:
-‘Here are the descriptions: {narration}.
-You are going to answer a multiple-choice question based on the descriptions.
-Here is the question: {question}?
-You must select one of these choices as the answer: A: {option_0} B: {option_1} C: {option_2} D: {option_3} E: {option_4} The correct answer is:’
-MovieChat [54].
-We use the official implementation of MovieChat (‘rese1f/MovieChat’, which incorporates LLaMa2-7B (‘meta-llama/Llama-2-7b’ and corresponding Video-LLaMa-7B (‘DAMO-NLP-SG/Video-LLaMA-2-7B-Finetuned’).
-Similarly, we use GPT-3.5 (‘gpt-3.5-turbo’) [42] to compute the accuracy of responses.
-Following the official implementation, we use the following prompt:
-‘By watching the video and combining it with the caption/subtitle from video , you are required to answer a question with multiple options: Question: {question}; Option A: {option_0}; Option B: {option_1}; Option C: {option_2}; Option D: {option_3}; Option E: {option_4}; Answer to the multiple-choice question based on the textual information extracted from the movie.
-If you don’t know the answer, please output ‘Unknown’.
-Your output should be just one of A, B, C, D, E and nothing else.’
-TimeChat [47].
-We use the official implementation of TimeChat (‘ShuhuaiRen/TimeChat-7b’).
-Following the official implementation, the prompt and the accuracy computation details are the same as those in the previous MovieChat model.
-For all models, we also incorporate subtitles as input.
-For that, we concatenate each query with the following text: ‘Subtitles: {subtitles}’.
-Subtitles are formatted as follows: ‘{start_timestamp} - {end_timestamp} {subtitle_i}’.
-We trim subtitles if they cannot fit into the input size.
-Results and analysis can be seen in Section 3.3 and Table 3.
-Appendix E Technical details on the temporal window study
-In this section, we present additional details and results for the temporal window study (see Section 3.2 and Figure 7).
-The goal of the temporal window study is to evaluate the performance of various methods on SFD at different temporal levels—shot, scene, and movie—across all modalities.
-Along with the discussed FrozenBiLM and LLoVi, we examine Video-LLaVA and LangRepo.
-At the shot-level, we truncate the input prompt to include only information from a specific shot.
-For FrozenBiLM and Video-LLaVA, it means that we only include subtitles and frames extracted from the shot.
-For LLoVi and LangRepo, we only include subtitles and captions extracted from the shot.
-Hence, we get logits for each answer option and shot, aggregated in a matrix of shape : we obtain a final prediction by taking the maximum logit over both dimensions.
-Note that the shot-level inference takes too much time to compute for LangRepo ( A100 GPU-hours per experiment), as this method involves many steps of rephrasing and summarizing.
-The scene-level inference is similar to the shot-level one, the main difference being that we constrain the input information (subtitles and frames/captions) to the scene-level, which arbitrarily corresponds to 10 shots in our case.
-We end up with a logit matrix of size .
-The movie-level inference is straightforward: it is the same as in Section D.
-Vision | Language | Vision-Language | ||
-FrozenBiLM | Shot-Level | 22.1 | 30.9 | 30.8 |
-Scene-Level | 22.7 | 34.8 | 34.8 | |
-Movie-Level | 23.4 | 38.2 | 38.6 | |
-Video-LLaVA | Shot-Level | 16.0 | 24.4 | 24.0 |
-Scene-Level | 23.7 | 20.1 | 20.1 | |
-Movie-Level | 34.2 | 21.3 | 24.7 | |
-LLoVi | Shot-Level | 32.8 | 38.5 | 50.1 |
-Scene-Level | 34.2 | 51.2 | 55.4 | |
-Movie-Level | 30.8 | 64.2 | 55.6 | |
-LangRepo | Shot-Level | - | - | - |
-Scene-Level | 26.9 | 27.9 | 28.1 | |
-Movie-Level | 29.0 | 32.1 | 31.0 | |
-Human | Movie-Level | 59.0 | 70.9 | 89.8 |
-The numerical results of the temporal window study are reported in Table 6.
-Along the temporal direction (i.e.
-row-wise for each method), we observe that: (i) generally, all methods exhibit consistent gains with larger temporal windows confirming the importance of long-term understanding for our designed questions, and (ii) exceptions are noted, such as with Video-LLaVa in Language and Vision-Language settings, likely due to its low precision (near random chance at 20%) which diminishes the impact of the temporal window.
-For different modalities (i.e.
-column-wise), we make the following observations: (i) Corroborating the argument in the main paper (Section 3.3), the Language-only models, in general, achieve better performances, e.g.
-LLoVi reaches the best performance at movie-level in the language-only setting (64.2%).
-(ii) The vision-only setting results in the weakest performance, revealing the largest gap compared to human performance (last row).
-This suggests significant potential for improvement in visual models.
-(iii) Current modality-fusion methods show limited effectiveness.
-VLM models like FrozenBiLM and Video-LLaVA only demonstrate marginal improvements when fusing vision and language, with performance increment from 38.2% to 38.6% for FrozenBiLM and from 21.3% to 24.7% for Video-LLaVa.
-However, all LLM-based methods experience a regression after modality fusion.
-Overall, the human benchmark highlights that there remains a substantial gap between the best methods and human performance, particularly in vision-language integration over extended temporal windows.
-Appendix F Human Study
-In this section, we present additional details for the conducted human study (see Section 3.3, ‘User Study’).
-We conducted a comprehensive human study to evaluate the upper-limit performance of our benchmark.
+These differences highlight the story-centric nature of the SF20K dataset.
+Subset statistics.
+In Table 7, we provide the key statistics of SF20K and its subsets.
+One key element to highlight is the annotation process.
+SF20K comes with almost 200K question-answer pairs, used for instruction tuning.
+Those have been generated automatically from subtitles.
+On the contrary, SF20K-Test questions have been automatically generated from human-written synopses, and subsequently manually curated.
+This applies also to SF20K-Test-Silent, which is simply a subset of SF20K-Test with only silent movies.
+Finally, SF20K-Test-Expert is a fully manual test set, where questions are designed by annotators who watched the movie.
+Question types.
+We use an LLM to automatically classify the question into 4 categories: setting-, character-, story-, and theme-related questions.
+In Table 10, we provide a few examples of questions for each question type from SF20K-Test.
+Appendix C Technical details on the data leakage experiment
+| Template: "In [MOVIE TITLE], [QUESTION]" |
+| In ’Titanic’, what happened to the Heart of the Ocean? |
+| In ’The Godfather’, how does Vito die? |
+| In ’Harry Potter and the Chamber of Secrets’, who is Dobby’s master? |
+| Question | Answer |
+| Setting | |
+| Where do John and Emma have their conversation? | On the bus. |
+| Which city is the backdrop? | Liverpool. |
+| What decade does the movie take place in? | 1990s. |
+| Character | |
+| Who is the main character? | Josh. |
+| What is the nationality of the soldier who becomes a prisoner? | German. |
+| How is Mr. Jones described? | As stern and harsh. |
+| Story | |
+| Can you name a specific tactic Tommy uses against Tiny? | He steals Tiny’s food at dinner and distracts her at a school-wide relay race. |
+| What is the imminent threat to Greg’s residence? | The house is in foreclosure and the bank is about to repossess it. |
+| What triggers Carla to believe her husband Frank will surprise her for their anniversary? | A bouquet left on her doorstep. |
+| Theme | |
+| What is a major theme? | The theme revolves around reconnecting and the complexities of honesty between exes. |
+| What main theme is explored through Sarah’s experience? | The theme of romantic expectations versus present realities. |
+| What is the main theme? | The processing of grief and reconciliation between siblings. |
+In this section, we include additional information on the data leakage study mentioned in Section 4.1.
+The objective of our data leakage study is to demonstrate that current large language models (LLMs) possess prior knowledge about commercial movies.
+This is because LLMs have been exposed to vast amounts of movie-related textual information during pretraining, including synopses, blog posts, news articles, and more.
+The experiment is set up as an open-ended question-answering task where the input consists of the movie title and corresponding question (see Table 9 for a few examples), and the output is a free-form text generated by the LLM.
+We evaluate the answer using an LLM to compute the similarity between the predicted and the correct answer, following [39].
+We choose ten of the most recent LLMs, among which six are open-source models: Gemma-2B, Gemma-7B, LLaMA-3-8B, LLaMA-3-70B, Mistral-7B, Mixtral-8x7B; and four commercial models: Claude 3 Haiku, Claude 3 Sonnet, GPT-3.5, GPT-4.
+The open-source models are run using the Hugging Face transformers library 555https://github.com/huggingface/transformers.
+The commercial models are run using their official API: the OpenAI API 666https://openai.com/api for GPT and the Anthropic API 777https://www.anthropic.com/api for Claude 3.
+We choose three movie datasets: MovieQA [61], LVU [70], and SF20K-Test (Ours).
+For MovieQA and SF20K-Test, the question-answer pairs are directly available.
+In the case of LVU, we re-formulate the classification tasks into templated questions to fit the open-ended question-answering format.
+For example, the director classification task is converted into the question ‘Who directed the movie [MOVIE TITLE]?’.
+The detailed results of the data leakage experiment are presented in Table 11.
+We notice that LLMs exhibit significantly higher data leakage on MovieQA and LVU datasets compared to SF20K.
+Models like GPT-4 and Claude 3 Sonnet answer correctly to more than 70% of the questions of both datasets, without any context except the movie title, and in an open-ended question-answering setup.
+Appendix D Human Study
+In this section, we present additional details for the conducted user study (see Section 4.2).
+We conducted a comprehensive user study to evaluate the upper-bound performance on our benchmark.
 Participants were asked to watch a movie and subsequently answer multiple-choice questions based on the content.
 To analyze the impact of different modalities, we implemented two ablation conditions: a muted version, where participants had access to the vision-only information without audio or subtitles, and a blind version, where participants were provided with the audio-only information, in the form of textual subtitles, without any video.
-To avoid any bias, we ensured that different sets of participants were assigned to the full, vision-only, and audio-only settings, preventing any overlap and potential influence from experiencing multiple conditions.
+To avoid bias, we ensured that different sets of participants were assigned to different settings (full, vision-only, and audio-only), preventing any overlap and potential influence from experiencing multiple conditions.
 Movies were randomly assigned to users to ensure a diverse range of data.
 The entire study was facilitated through an integrated interface, which restricted access to external information such as movie loglines and synopses, with only the movie title being displayed (see Figure 9).
-Appendix G Technical details on data leakage experiment
-Name | SFD | MovieQA | LVU |
+| Name | SF20K-Test | MovieQA | LVU |
 | Open-source models | |||
@@
 | GPT-4 | 31.5 | 71.3 | 76.0 |
-In this section, we include additional information on the data leakage study mentioned in the main paper Section 3.1.
-The objective of our data leakage study is to demonstrate that current large language models (LLMs) possess prior knowledge about commercial movies.
-This is because LLMs have been exposed to vast amounts of movie-related textual information during pretraining, including synopses, blog posts, news articles, and more.
-The experiment is set up as an open-ended question-answering task where the input consists of a movie title and a corresponding question, and the output is a free-form text generated by the LLM.
-We then evaluate the answer using an LLM to compute the similarity between the predicted and the correct answer, following [37].
-We choose ten of the most recent LLMs, among which six are open-source models: Gemma 2B, Gemma 7B, LLaMA-3 8B, LLaMA-3 70B, Mistral 7B, Mixtral 8x7B; and four commercial models: Claude 3 Haiku, Claude 3 Sonnet, GPT-3.5, GPT-4.
-The open-source models are run using the Hugging Face transformers library 666https://github.com/huggingface/transformers.
-The commercial models are run using their official API: the OpenAI API 777https://openai.com/api for GPT and the Anthropic API 888https://www.anthropic.com/api for Claude 3.
-We choose three movie datasets: MovieQA [56], LVU [64], and SFD (ours).
-For MovieQA and SFD, the question-answer pairs are directly available.
-In the case of LVU, we re-formulated the classification tasks into templated questions to fit the open-ended question-answering format.
-For example, the director classification task is converted into the question ‘Who directed the movie {movie_title}?’.
-The detailed results of the data leakage experiment are presented in Table 7 and Figure 10.
-Their analysis shows that LLMs exhibit significantly higher data leakage on MovieQA and LVU datasets compared to SFD.
-Models like GPT-4 and Claude 3 Sonnet answer correctly to more than 70% of the questions of both datasets, without any context except the movie title.
-The scatter plot emphasizes a correlation between data leakage and model knowledge (i.e.
-MMLU performance), with higher MMLU scores correlating with increased leakage issues on compromised datasets.
-This trend is consistent across both open-source and commercial models, highlighting the vulnerability of larger and more advanced models to memorizing training data.
-In sharp contrast, the relatively low and flat performance trend on SFD suggests it is a more robust benchmark for evaluating model capabilities.
-These findings emphasize the need for robust evaluation datasets to accurately assess model performance and mitigate the impact of data leakage.
-Appendix H Samples from SFD
-In the following pages, we show some samples from SFD.
+| Video ID | Question | Answer | Prediction |
+| Visual cues | |||
+| IehPbS9fbPY | How does Ernie’s suicide attempt fail?
+| The attempt is unsuccessful and results in a hole in his ceiling.
+| Ernie’s suicide attempt fails when he is interrupted by a significant moment of connection or realization, leading him to reconsider his actions.
+|
+| xErsPPWDDdk | How do ordinary objects appear to Yumi during her visions? | Ordinary objects appear as ladles to Yumi during her visions. | I’m unable to provide specific details … |
+| Character identification | |||
+| wD–FpKwnvA | Who does Corey live with? | She lives with her grown daughter. | Corey lives with his mother, Karen. |
+| mhxEnIp7YIg | Who invites Jane to the boxing gym? | Tisha | I’m unable to identify … |
+| Temporal understanding | |||
+| 0PKtGVJEpdg | Where does Becca encounter the man again after the incident? | She meets him at a bar one evening. | Becca encounters the man again after the incident in an alley. |
+| 3PejXWtS4xs | What is Harris doing in the apartment at the beginning of the movie?
+| Waiting for his target to arrive.
+| At the beginning of the movie, Harris is in the apartment, having a conversation on the phone.
+|
+Appendix E Technical details on baselines
+In this section, we present additional details for the baselines discussed in Section 4.3, where we report the benchmark performance of the state-of-the-art video understanding methods on SF20K-Test.
+We detail the models evaluated, the prompting strategy, and the metrics used.
+Models.
+We evaluate eleven open-source models: FrozenBiLM [6], mPLUG-Owl2 [80], TimeChat [52], Video-Llava [36], MovieChat [59], Llava-OneVision [32], Long-Llava [67], LongVA [84], Llava-Next-Video [85], Llama-3.2-Vision [43], and Pixtral [3]; and two commercial models: LloVi [83] (based on GPT-3.5 [45]), and GPT-4o-mini [46].
+Most open-source models are available on Hugging Face’s transformers library 888https://huggingface.co/docs/transformers, except for FrozenBiLM 999https://github.com/antoyang/FrozenBiLM, mPLUG-Owl2 101010https://github.com/X-PLUG/mPLUG-Owl, and LongVA-7B 111111https://github.com/EvolvingLMMs-Lab/LongVA.
+Commercial models are available through their official API.
+Note that commercial models are not free; therefore, we are limited in their usage.
+We are only able to evaluate GPT-4o-mini on 10% of SF20K-Test (i.e.
+approximately 500 questions).
+Prompting. We prompt all the models with the same queries, inspired by Gemini [51]. For multiple-choice question answering, we use the following prompt:
+‘You will be given a question about a movie.
+Try to answer it based on the subtitles and the frames from the movie.
+Question: [QUESTION] Possible answer choices: (1) [OPTION 1] (2) [OPTION 2] (3) [OPTION 3] (4) [OPTION 4] (5) [OPTION 5] Output the final answer in the format "(X)" where X is the correct digit choice.
+DO NOT OUTPUT with the full answer text or any other words.‘
+For open-ended question answering, we use the following prompt:
+‘You will be given a question about a movie. Try to answer it based on the subtitles and the frames from the movie. Question: [QUESTION] Answer it shortly and directly without repeating the question.‘
+We include audio information by adding the transcript in the prompt, in the following format: ‘[START TIMESTAMP] - [END TIMESTAMP]: [SUBTITLE]‘.
+In case of modality ablation, we simply remove either the subtitles or the video frames from the inputs and adapt the query accordingly.
+| Multiple-Choice QA | ||||||
+| V | L | VL | ||||
+| FrozenBiLM | ||||||
+| Shot-level | 22.1 | 30.9 | 30.8 | |||
+| Scene-level | 22.7 | +0.6 | 34.8 | +3.9 | 34.8 | +4.0 |
+| Movie-level | 23.4 | +1.3 | 38.2 | +7.3 | 38.6 | +7.8 |
+| Video-Llava-7B | ||||||
+| Shot-level | 16.0 | 24.4 | 24.0 | |||
+| Scene-level | 23.7 | +7.7 | 20.1 | -4.3 | 20.1 | -3.9 |
+| Movie-level | 34.2 | +18.2 | 21.3 | -3.1 | 24.7 | +0.7 |
+| Llovi | ||||||
+| Shot-level | 32.8 | 38.5 | 50.1 | |||
+| Scene-level | 34.2 | +1.4 | 51.2 | +12.7 | 55.4 | +5.3 |
+| Movie-level | 30.8 | -2.0 | 64.2 | +25.7 | 55.6 | +5.5 |
+| Human | ||||||
+| Movie-level | 59.0 | 70.9 | 89.8 | |||
+Metrics.
+For multiple-choice QA, we use regular expression to extract the digit choice inside the parenthesis in the model’s response.
+We then compute the overall accuracy.
+For open-ended QA, we use a LLM to compare the free-form response from the model to the correct answer.
+In details, we prompt GPT-3.5 (‘gpt-3.5-turbo-0125‘) for the OpenAI API, and use the prompt provided by [39].
+We get a yes/no value and an associated score.
+We then compute the average score and the overall accuracy.
+Appendix F Failure cases analysis
+In this section, we provide an analysis of failure cases to better understand the limitations of our model in video question answering.
+Table 12 categorizes these failure cases into three key domains: visual cues, where the model struggles to interpret visual details critical to answer the question; character identification, where the model fails to correctly associate characters with their names, actions, or attributes; and temporal understanding, which highlights challenges in reasoning about events over time.
+By examining these errors, we aim to emphasize specific weaknesses and inform future improvements in multimodal reasoning and long-term video understanding.
+Appendix G Technical details on the temporal window study
+In this section, we present additional details and results for the temporal window study (see Section 4.5).
+The goal of this experiment is to evaluate the performance of various methods on SF20K-Test at different temporal levels—shot, scene, and movie—across all modalities.
+We examine FrozenBiLM [6], LLoVi [83], and Video-Llava-7B [36].
+At the shot-level, we truncate the input prompt to include only information from a specific shot.
+For FrozenBiLM and Video-LLaVA, it means that we only include subtitles and frames extracted from the shot.
+For LLoVi and LangRepo, we only include subtitles and captions extracted from the shot.
+Hence, we get logits for each answer option and shot, aggregated in a matrix of shape : we obtain a final prediction by taking the maximum logit over both dimensions.
+The scene-level inference is similar to the shot-level one, the main difference being that we constrain the input information (subtitles and frames/captions) to the scene-level, which arbitrarily corresponds to 10 shots in our case.
+We end up with a logit matrix of size .
+The movie-level inference is straightforward: it is the same as in Section 4.3.
+The numerical results of the temporal window study are reported in Table 13.
+Along the temporal direction (i.e.
+row-wise for each method), we observe that: (i) generally, all methods exhibit consistent gains with larger temporal windows confirming the importance of long-term understanding for our designed questions, and (ii) exceptions are noted, such as with Video-Llava-7B in Language and Vision-Language settings, likely due to its low precision (near random chance at 20%) which diminishes the impact of the temporal window.
+Appendix H Prompts
+We provide the prompts for the data generation process: question generation prompt (see Table 14), distractor generation prompt (see Table 15), and question classification prompt (see Table 16).
+As detailed in Section 3.2, those prompts were used with GPT-4 (’gpt-4-turbo-2024-04-09’) through the OpenAI API.
+![[Uncaptioned image]](x13.png)
+![[Uncaptioned image]](x14.png)
+![[Uncaptioned image]](x15.png)
+Appendix I Samples from SF20K-Test
+In the following pages, we show some samples from SF20K-Test.
 For each sample, we display the title, the logline, the synopsis, a few frames, some subtitles, and the designed multiple-choice questions.
-See Figures 11, 12,13, 14.
-Appendix I Prompts
-We provide the prompts for the data generation process: question generation prompt (see Figure 8), distractor generation prompt (see Figure 9), and question classification prompt (see Figure 10).
-As detailed in Section 2.2 , those prompts were used with GPT-4 (’gpt-4-turbo-2024-04-09’) through the OpenAI API.
-![[Uncaptioned image]](x20.png)
-![[Uncaptioned image]](x21.png)
-![[Uncaptioned image]](x22.png)
+See Figures 10, 11,12, 13.
 Appendix J Datasheet for Datasets
@@
 For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
-This dataset and benchmark are created to assess long-form video understanding capabilities of modern multimodal systems.
-It is different from other video question-answering datasets as it features longer videos and associated story-related questions.
-Additionally, it stands out as a dataset that is available online and not exposed to current large language models (limited data leakage issues).
+This dataset is created to improve long-form video understanding capabilities of modern multimodal systems.
+It comes with a benchmark to assess current performance.
+This video question-answering dataset differs from previous ones as it features longer videos and associated story-related questions.
+Additionally, it stands out as a dataset that is available online and not exposed to current large language models (which limits data contamination issues).
 Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
-This dataset was created by Ridouane Ghermi††footnotemark: , Xi Wang††footnotemark: , Vicky Kalogeiton††footnotemark: , and Ivan Laptev999MBZUAI.
+This dataset was created by Ridouane Ghermi††footnotemark: , Xi Wang††footnotemark: , Vicky Kalogeiton††footnotemark: , and Ivan Laptev121212MBZUAI.
 Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
@@
 Please provide a description.
-Our dataset instance consists of a short movie (around 5-20 mins), some metadata (e.g.
-movie title, logline, synopsis, country of origin, language, release year), and a list of multiple-choice questions and answers.
+A single instance consists of a short movie (around 5-50 mins), and corresponding metadata (i.e. movie title, logline, country of origin, language, release year).
 How many instances are there in total (of each type, if appropriate)?
-There are 1,078 movies and 4,885 questions. Each question comprises one correct answer and four wrong answers.
+There are 20,143 movies.
 Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
@@
 If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
-The dataset is not a sample of instances from a larger set.
+The dataset is new, this is not a sample from a larger set.
 What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
-Each instance includes a movie as a video file. Metadata comes as .csv files, in the form of text. We also provide processed features.
+Each instance consists of a movie (video file), metadata (csv file), and extracted annotations (i.e. shot boundaries, subtitles, captions, and face tracks)
 Is there a label or target associated with each instance? If so, please provide a description.
@@
 Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
-We provide a dataframe that links question ids and video ids.
+Not applicable.
 Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
-SFD is a test set only.
+SF20K contains a train set and a test set, called SF20K-Test. Subsequently, SF20K-Test contains two subsets: SF20K-Test-Silent and SF20K-Test-Expert.
 Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
-No, the dataset has been manually curated.
+No.
 Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
@@
 Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
-All movies are available on YouTube; We only provide the corresponding URLs.
-It should remain constant over time as it comes from YouTube channels that curate a movie catalogue.
+All movies are available on YouTube and Vimeo.
+We only provide URLs.
+It should remain constant over time as it comes from YouTube and Vimeo channels that curate a movie catalogue.
 The dataset is distributed under the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license.
@@
 If so, please describe how.
-Movies are available online and downloaded from YouTube using yt-dlp, along with metadata. Multiple-choice questions are generated by a commercial LLM, GPT-4, and further manually curated.
+Movies and metadata are available online, and downloaded directly from YouTube and Vimeo.
 What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)?
 How were these mechanisms or procedures validated?
-Video URLs are scraped using the official YouTube API.
-Videos and metadata are downloaded using yt-dlp.
-Questions are generated using GPT-4 through the official OpenAI API and further refined by manual curation.
+Video URLs are scraped using the official APIs. Videos and metadata are downloaded using yt-dlp.
 If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
@@
 If not, you may skip the remainder of the questions in this section.
-We extracted several annotations from the dataset, see Sections C.
+We extracted several annotations from the dataset, see Sections B.
 Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
@@
 Has the dataset been used for any tasks already? If so, please provide a description.
-The dataset is accompanied by two tasks: multiple-choice and open-ended question answering.
+The dataset is used to train video question answering models.
 Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
-Yes, we will make this available on the dataset GitHub repository.
+Yes.
 What (other) tasks could the dataset be used for?
@@
 We provide video URLs and annotations.
-Additionally, we provide code to download videos, pre-process the data and run baselines.
+Additionally, we provide code to download videos, pre-process the data, run baselines, and fine-tune models.
 When will the dataset be distributed?
-The dataset is already available.
+After acceptance.
 Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
@@
 If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
-The dataset is a test benchmark, it will remain stable to ensure fair evaluation of future methods.
+The dataset might be updated to increase its size and diversity. However, the test set will remain stable to ensure fair evaluation of future methods.
 If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
@@
 If so, please provide a description.
-If others want to contribute to the dataset, by adding more tasks for example, we would gladly host them in our repository.
+If others want to contribute to the dataset, by adding videos or labels for example, we would gladly host them in our repository.
 Also, we will provide as many details and tools as possible to reproduce our annotation process and create more tasks.
 Any other comments?
-N/A.